A Complete Bibliography of Applied Statistics: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

12 August 2014 Version 4.08

Title word cross-reference

2 × 2 [372, 595]. 2 × K [401]. χ^{2} [245, 122]. F [566]. J [575]. k [88]. κ [380]. L_{1} [265, 94]. L_{∞} [94]. N [575, 193, 368, 418]. p [504, 296]. QR [174]. t [324, 247, 562]. W [474].

-coefficient [380]. -distributed [562]. -distribution [566]. -Norm [265]. -test [474]. -Tests [324]. -Values [504, 296].

1 [145]. 1.1 [433]. 1.10 [172]. 1.2 [404]. 1.25 [215]. 1.3 [115]. 109 [59, 123, 124, 4]. 121 [175]. 139 [340]. 152 [147]. 179 [543]. 181 [474]. 185 [148]. 189 [473]. 1986 [18]. 199 [149].

2.05 [215]. 206 [125, 126]. 221 [35, 34]. 226 [60, 475, 5]. 229 [388]. 238 [265]. 251 [353]. 252 [31]. 253 [32]. 254 [56]. 255 [57]. 256 [58]. 257 [80]. 258 [81]. 259 [82]. 260 [117]. 261 [118]. 262 [119]. 263 [120]. 264 [121]. 265 [145]. 266 [146]. 267 [173]. 268 [174, 490]. 269 [216, 299]. 270 [217]. 271 [218, 266]. 272 [219]. 273 [243]. 274 [244, 415, 416]. 275 [245]. 276 [246]. 277 [263]. 278 [264]. 279 [504, 296]. 280 [297]. 281 [298]. 282 [323]. 283 [324]. 284 [338]. 285 [339]. 286 [351]. 287 [352]. 288 [371]. 289 [372]. 290 [373]. 291 [387]. 292 [400]. 293 [401]. 294 [402]. 295 [414]. 296 [445]. 297 [446]. 298 [471]. 299 [472]. 2nd [73, 46, 26, 204, 235, 318, 312]. 3.0 [144]. 300 [489]. 301 [502]. 302 [503].

303 [514]. 304 [515]. 305 [526]. 306 [527]. 307 [528]. 308 [538]. 309 [539]. 310 [540]. 311 [541]. 312 [542]. 313 [553]. 314 [554]. 315 [562]. 316 [563]. 317 [565]. 318 [566]. 319 [574]. 320 [575]. 321 [576]. 3rd [320, 205].

4 [417]. 4th [103, 213, 69].

5V [116].

6.01 [55]. 63 [4, 1]. 64 [4, 2]. 65 [33, 3].

76 [247].

91 [122].

A. [111, 27, 316, 25, 259, 258, 202, 234, 208, 23, 200, 107, 21]. Aalen [155]. ABACUS [391]. ability [511]. absolute [309]. Accelerated [209, 582]. acceptance [191]. accident [87]. Accounting [132]. accuracies [589]. Acid [250, 387]. Acknowledgement [649]. acquired [610, 494, 423]. Adams [18]. Adaptive [226, 466, 577, 352]. added [271]. Addition [162]. Additive [203]. Additivity [338]. adequacy [383]. adjusting [273]. adjustment [42]. Advanced [49]. Advances [70]. Against [502]. Age [361, 657]. ages [42]. aggregates [156]. Agricultural [482, 522, 182]. air [643]. airline [596]. Aitkin [19]. Akaike [139]. Albert [259]. Albeverio [20]. Algorithm [471, 149, 553, 414, 542, 562, 453, 566, 565, 446, 324, 32, 147, 59, 123, 124, 338, 471, 82, 217, 542, 472, 540, 298, 297, 4, 502, 35, 34, 219, 117, 118, 245, 553, 218, 246, 266, 489, 353, 400, 58, 265, 81, 175, 60, 526, 575, 543, 80, 125, 148, 126, 402, 387, 247, 145, 445, 323, 576, 372, 401, 539, 56, 504, 388, 296, 574, 31, 475, 264, 216, 299, 5, 339, 173, 514, 1, 2, 119, 57, 541, 562, 243, 244, 415, 373, 414, 416, 527, 563, 146, 263, 371, 33, 120, 121, 554, 351, 515].

Algorithm [3, 528, 474, 340, 122, 503, 174, 473, 490, 566, 538, 352]. Algorithms [36, 580, 544, 555, 567, 446, 324, 32, 147, 59, 124, 338, 471, 82, 217, 472, 149, 298, 297, 4, 502, 35, 219, 117, 118, 245, 218, 246, 266, 489, 353, 400, 58, 265, 81, 175, 60, 526, 80, 148, 126, 402, 387, 247, 145, 445, 323, 372, 401, 56, 504, 388, 296, 31, 475, 264, 216, 299, 5, 339, 173, 514, 1, 2, 119, 57, 243, 244, 373, 414, 416, 527, 146, 263, 371, 33, 120, 121, 351, 515, 3, 528, 474, 340, 122, 503, 174, 473, 490, 352]. allocation [182]. Allowing [659]. Alphabetic [387]. Alternative [253, 395]. Alwan [451]. Among [510]. Analyse [206]. analysed [644]. Analyses [280, 334, 209, 239]. Analysing [648, 439, 168, 511, 497]. Analysis [223, 310, 347, 470, 14, 103, 285, 184, 27, 523, 559, 358, 558, 346, 199, 259, 253, 655, 361, 207, 312, 430, 590, 277, 107, 102, 431, 229, 510, 170, 69, 79, 573, 606, 114, 52, 139, 201, 39, 467, 483, 465, 654, 607, 520, 508, 42, 653, 231, 509, 533, 587, 638, 557, 187, 131, 381, 602, 12, 7, 333, 532, 468, 630, 562, 613, 225, 136, 43, 453, 424, 494, 659, 572, 179, 641, 616, 366, 379, 46, 165, 169, 234, 48, 53]. Anderson [19, 21]. animal [362]. Annual [329, 328]. applicants [639]. Application [520, 617, 357, 162, 568, 534, 444, 425, 252, 484, 228, 612, 662, 618, 657, 408, 602, 13, 656, 625, 438, 423, 628, 91, 571]. Applications [470, 202, 430, 453, 159, 652, 207, 45, 110, 291, 258, 212, 197, 72, 238]. Applied [236, 54, 239, 140, 653, 182, 613, 616, 144, 519]. Approach [487, 281, 455, 456, 397, 92, 367, 408, 615, 302, 153, 636, 443]. approaches [605]. Appropriate [485]. Appropriateness [512]. Approximating [181]. Approximation [44, 94]. aquaculture [637]. Arbitrary [371]. archaeological [622]. Archaeometry [470]. area [538]. arising [611]. arithmetic [554]. arrays [636]. Arrest [510]. Arthur [207]. aspects [587]. Assessing [455, 570, 252, 604, 564, 569]. assessment [582]. Assignments [140]. Association [32, 551, 511]. Assurance [237]. Asymptotic [524]. Athletics [469, 537, 654]. attacked [331]. attained [436]. attributable [616]. Attribute [333]. Australian [306]. Author [579]. authors [661]. Autocorrelated [127]. Autocorrelations [504, 296]. Automatic [148]. Automation [47]. autoradiographic [465]. autoregressive [541, 89]. autoregressive-type [89]. Average [81, 526, 541, 594]. Averaged [385]. averages [627]. averaging [569]. Axes [310].

B [167, 74, 53, 21, 79, 29]. B. [24, 290, 23, 105, 19, 106, 114]. Back [37, 61, 83, 128, 151, 176, 221, 248, 267, 300, 325, 341, 354, 377, 392, 406, 419, 434, 447, 461, 476, 491, 505, 517, 529]. Backcalculation [133]. bacteriological [411]. Bain [312, 163]. Balanced [527]. bark [331]. bark-beetles [331]. Barkman [101]. Barrett [164]. base [276]. base-line [276]. based [600, 648, 563]. basis [193, 636]. Batey [70]. Bayes [261, 622, 502, 288, 134, 254]. Bayesian [610, 455, 231, 523, 327, 599, 619, 90, 220, 408, 602, 562, 428, 644, 153, 588, 364, 569, 616, 366, 379, 368, 418, 108, 443]. Becker [102]. beetles [331]. before [202]. behaviour [631]. Belief [24, 169]. Belsley [257]. Bench [67]. Bench-Marking [67]. Bench-Marks [67]. Best [173, 404]. Beta [59, 124, 540, 4, 60, 512, 475, 5, 1, 2, 473, 123]. Beta-Binomial [512]. better [135]. betting [551]. between [484, 41, 308]. Bhattacharya [197]. biadditive [524]. Biemer [50]. Binary [46, 314, 196, 359, 520, 653, 189, 426, 615, 630]. Binomial [512, 473, 368, 418, 584, 193]. biogeography [408]. Biological [319, 493]. biomedical [533, 618]. Biopharmaceutical [28]. Birth [443, 617]. birthweight [42]. Bit [121]. Bivariate [291, 292, 247, 373, 263, 528, 576, 538, 330, 274]. Blanchard [20]. Block [27, 527]. BMDP [262, 116]. Board [487]. bone [511]. Book [198, 319, 71, 315, 167, 313, 18, 103, 111, 73, 27, 70, 110, 320, 260, 291, 292, 316, 74, 25, 46, 77, 199, 259, 165, 26, 258, 24, 212, 113, 204, 169, 206, 235, 202, 210, 317, 314, 293, 211, 234, 236, 48, 207, 208, 205, 318, 321, 290, 312, 49, 68, 29, 23, 28, 45, 76, 164, 109, 197, 209, 72, 200, 107, 102, 203, 54, 168, 239, 53, 257, 240, 214, 75, 213, 170, 104, 163, 105, 47, 101, 69, 19, 112, 141, 21, 140, 106, 22, 51, 114, 52, 20, 166, 50, 139, 237, 108, 261]. Book [238, 201]. bootstrap [425, 625]. bootstraps [179]. borehole [382]. both [132]. Bound [149]. Bovine [557, 559, 558, 556]. Box [219, 162, 253, 8]. Box-Jenkins [162]. Bradley [444]. Bradley-Terry-Type [444]. brain [532, 586]. Branch [149]. branching [635]. Breakdown [323]. breast [520]. Britain [568, 497]. buffer [583]. Bullen [22]. Bunday [290]. Bunke [68]. bus [578, 549]. Business [26].

C [318, 23, 197]. C.

[167, 103, 291, 292, 170, 69]. Calculating [247]. Calculo [320]. Calculus [320]. Calibration [332, 41, 611, 141]. Canadian [593]. cancer [91]. cancers [520]. canonical [629]. Capability [64, 275]. Cardiac [510]. Carlo [617, 192, 547, 534, 497, 572]. carry [603]. carry-over [603]. Case [487, 253, 413, 510, 161, 559, 593, 588, 538]. Case-Control [510, 559]. Case-Study [487, 413, 593, 588]. cases [328]. Catch [361]. Catch-At-Age [361]. Categorical [279, 460, 440, 114, 427, 607, 131, 439]. categories [135, 160]. categorized [306]. category [521]. Cattle [14]. Caulcutt [451]. Causation [24]. cause [584]. caused [455]. CD4 [571, 494]. cell [571]. cells [602].

Censored [400, 340, 381, 7, 423]. censoring [284]. Census [334]. Central [245, 60, 247, 475, 540, 566]. centrality [566]. cervical [91]. Ch. [206]. chain [626, 617, 547, 534, 639, 572]. Chains [507, 365]. Change [24]. changed [650]. changepoint [231, 43, 364, 411]. changepoints [328]. Changes [512]. Chaos [315]. Characterizations [27]. Charpy [572]. Charts [450, 526, 594, 191]. Chatfield [103, 69]. checking [179]. Chemical [633]. chess [13]. Chhikara [45]. children [597, 657]. choice [11]. CIA [79]. Cinlar [313]. Circuit [583, 487]. class [568, 618, 306]. Classes [31]. classical [41]. Classification [218, 246, 266, 633, 369, 599, 619]. clearance [625]. Clinical [259, 189]. Cluster [165, 431, 587]. Clustering [135, 386]. clusters [439]. coastal [605]. Codes [334]. Coefficient [117, 118, 380]. Coefficients [298, 458]. Cohen [23]. cohort [559, 558, 556, 557, 367]. Colenutt [164]. Collapsibility [402]. Collett [314]. Collinearity [257]. Column [482, 546, 96]. Columns [57]. combination [224]. Combinatoric [218, 246, 266]. Combined [86, 487]. Combining [592]. comments [154, 568]. Common [106, 592, 584]. Companion [321]. comparative [508, 596]. Comparing [150, 243, 625, 589, 656, 652]. Comparison [223, 486, 605, 157, 647, 646, 413, 10, 166, 612, 662, 131, 13, 93]. Comparisons [232, 601]. competing [276]. complete [484]. Completion [344]. Complex [170]. compliance [560]. complicated [96]. Component [470, 613, 136, 616]. Components [106, 286]. Compositional [251, 345]. Compositions [345]. Compounds [633]. Computation [324, 502]. computer [583, 135]. Computing [540, 245, 60, 372, 401, 504, 388, 296, 475, 5, 240, 566]. Concave [352]. concentration [536].

Concepts [391]. concomitant [564, 604]. concrete [92]. conditional [511, 436, 158]. Conditioning [257]. Confidence [82, 272, 373, 79, 524, 192, 538, 396]. Confined [340]. Confocal [602]. Congdon [70]. Conjointe [206]. Conrad [140]. consequences [479, 210]. constancy [384]. constrained [656]. Constructing [153]. Construction [553, 120, 546]. constructs [581]. Containing [173]. Contaminant [464]. Content [398]. Contents [620, 660, 554]. Contingency [502, 148, 402, 436, 158, 160]. Continuation [385]. Continuous [291, 292, 381, 394]. Contrasts [570]. Control [449, 86, 450, 162, 526, 317, 318, 510, 101, 139, 183, 559, 302, 594, 191, 75]. controlled [560]. Controlling [629]. controls [362]. conversion [93]. Converting [156]. CORANA [294]. Cornell [198]. Cornish [216, 299]. Cornish-Fisher [216, 299]. Corporation [78]. Correcting [280]. Correction [123, 124, 149, 35, 34, 266, 353, 543, 125, 126, 299, 415, 416, 490, 9]. corrections [658]. Correlated [409, 359, 520]. Correlation [117, 118, 353]. Correspondence [229, 333]. Corrigendum [619, 662, 220, 577, 150, 578, 460, 516, 127, 418]. Count [595, 467, 308, 481]. counting [441]. Counts [15, 329, 571, 536, 494, 521, 91]. Course [211, 507, 21, 71]. Covariance [430, 252]. covariates [612, 662, 394, 443]. Cox [260, 46, 253, 8, 63]. CP [480]. Cramer [274]. Cramer-von [274]. creatinine [625]. credible [538]. credit [639]. Creed [519]. Crime [250, 308]. criteria [564, 604]. critical [485]. crop [96]. Cross [57, 438, 88]. cross-sectional [438, 88]. Cross-Term [57]. Crossover [553, 603, 427, 189, 595]. Crowder [199]. Cumulative [86, 147, 157, 81, 526, 183, 410]. Currents [545]. Curvature [396]. curve [410, 133]. Curves [486, 548, 272, 494]. Cyrus [433].

Cytel [78].

D [27, 291, 292, 258, 207, 290, 168, 104, 79, 238, 414]. D-Optimal [414]. D. [260, 74, 46, 77, 199, 113, 314, 318, 404, 164, 257, 170, 19, 22, 52, 238]. Daintith [104]. Data [223, 251, 285, 523, 344, 46, 358, 346, 400, 190, 165, 90, 220, 402, 317, 548, 56, 430, 209, 119, 460, 440, 395, 257, 428, 345, 340, 334, 21, 573, 130, 386, 114, 52, 166, 409, 359, 467, 483, 66, 427, 484, 607, 520, 156, 597, 189, 638, 596, 426, 560, 655, 367, 187, 131, 381, 308, 12, 7, 584, 481, 593, 578, 549, 571, 630, 595, 534, 601, 550, 613, 631, 453, 424, 438, 644, 88, 659, 594, 572, 606, 423, 641, 382, 284, 314, 102, 168, 114]. DATACHAIN [295]. dating [614]. David [260, 55, 349]. dead [452]. decay [638]. December [182]. Decennial [334]. Decision [24, 172, 169, 48, 29, 522, 153]. decision-making [522]. Decisions [172]. Decomposability [402]. Decomposing [575]. Decomposition [490, 425, 174]. defects [443]. deficiency [610, 494, 423]. degrees [562]. Delaney [168]. dementia [589]. Demography [70]. Denes [258]. Deniz [171]. Densities [217]. Density [185, 352, 493, 538, 561]. Deoxyribonucleic [250, 387]. Department [417]. departures [511, 273]. dependence [63]. Depth [528]. Design [583, 653, 487, 357, 414, 201, 508, 587, 225, 215]. DESIGN-EASE [215]. DESIGN-EXPERT [215]. designed [641]. Designing [168]. Designs [454, 194, 437, 281, 482, 546, 656, 603, 286, 96, 652, 27]. Detecting [650, 180, 568, 411]. detection [464, 651, 271, 283]. detector [637]. deterioration [92]. determinations [397]. Determining [149]. Development [28, 112]. Developmental [409]. Developments [258]. Deviance [184]. devising [96]. DEX [171]. diabetes [564, 604]. Diagnosability [39]. diagnostic [564, 604]. Diagnostics [257]. Diagrams [169]. Dictionary [104]. differ [442, 272]. Difference [652, 41]. differenced [382]. Differences [173]. Differential [280]. Diffusion [539]. Digby [71]. dilution [536]. Dimensional [343, 463, 347]. direct [558]. Directional [363, 114, 483]. Dirichlet [146]. Disaster [287]. discordant [559]. discovering [456]. Discrete [535, 385, 626, 94, 639]. Discriminant [430]. discrimination [561]. Discussion [451, 610, 641]. Disease [102, 507, 252, 615, 134]. diseases [585, 640]. Dispersion [483, 136, 273, 652]. dissolution [509]. distance [655, 485]. distributed [562]. Distribution [18, 338, 498, 540, 35, 117, 118, 245, 58, 512, 387, 264, 45, 146, 473, 34, 365, 122, 88, 566]. Distributions [291, 400, 372, 401, 371, 516, 565, 9, 7, 648, 640, 429, 423, 616, 292]. DNA [650, 626]. DO [575]. Dobson [200]. domain [643]. Dose [285, 409, 442]. Dose-Response [285, 409]. Dotplots [499]. double [443]. doubly [284]. Down [617]. Driessen [47]. Drop [358, 571, 630, 645]. Drop-Out [358, 630, 645]. drop-outs [571]. Drug [28, 112, 153, 224]. Dual [388]. Durbin [504, 296]. Durbin-Watson [504, 296]. DYMO [241]. Dynamic [90, 220, 646, 240, 139, 411]. Dynamically [575].

E.

[319, 313, 260, 46, 259, 235, 197, 239, 214, 47]. Early [194, 42]. Earnings [18]. EASE [215]. Ecologic [280]. ecological [479]. Economic [26]. Edge [386]. Editors [138, 525, 100, 233, 289, 270, 422, 370, 17, 256, 552, 501, 335, 432, 513, 311, 459, 399, 608, 609]. Edn [46, 73, 320, 235, 318, 312, 103, 26, 204, 205, 213, 69]. education [568]. Edwards [349]. Effect [632, 512, 412, 612, 662, 560, 586, 645]. effectiveness [634]. Effects [232, 647, 189, 618, 603, 160]. efficacy [495]. efficiency [645]. Efficient [489, 546, 566]. Elementary [76, 213]. emission [611]. Emphasizing [291]. Empirical [622, 166, 154, 134, 261]. Employer [77]. Employment [51]. encephalopathy [559, 558, 556, 557]. Encyclopedia [167]. end [9]. energy [572]. Engelhardt [312, 163]. Engineer [75]. Engineering [292, 237, 20]. Enumeration [543]. Environment [240, 524]. Envisioning [214]. epidemic [610]. epidemiology [547]. EPILOG [143]. equation [367]. Error [351, 507, 592]. Error-in-Variables [351]. errors [562, 309, 625]. errors-in-variables [625]. estimate [493]. estimates [380]. Estimating [82, 74, 193, 536, 438, 507, 88, 186, 521, 329, 605, 367, 410, 495, 436, 286, 616]. Estimation [32, 457, 35, 632, 282, 323, 539, 56, 23, 146, 97, 254, 306, 365, 351, 340, 473, 252, 166, 368, 418, 565, 622, 34, 657, 381, 182, 159, 611, 423, 412, 636, 561, 646]. Estimator [196, 460, 440]. Estimators [446, 195, 134, 10]. Evaluating [281]. Evaluation [522, 117, 164, 161]. events [468]. Everitt [105]. Every [75]. Evidence [558, 161]. evolution [379]. Exact [297, 372, 401, 601, 436, 613, 158, 371, 576, 541, 497]. Example [114]. Exceptional [469, 537]. excess [11]. Exchange [414, 511, 166]. exchange-ability [511]. exciting [89]. Exemplified [387]. exhibit [638]. Exogenous [539]. Expansion [338, 216, 299]. experiment [452, 656]. Experimental [194]. Experiments [198, 255, 303, 168, 437, 65, 583, 427, 535, 96, 645, 641, 646, 52, 201]. expert [479, 215]. Explanatory [458, 550]. Exploration [25]. Exploratory [557, 7]. exploring [647]. exponentially [594]. Expression [130]. Extended [13]. extensions [655]. extra [11]. extra-Poisson [11]. Extrabinomial [285]. Extraneous [460, 440]. extravariation [12]. Extremal [654].

Extreme [363, 523, 632, 346, 400, 545, 186]. Extremes [357].

F [110, 55, 204, 170, 47]. facilitate [225]. Factor [502, 166]. Factorial [454, 482, 483, 607]. Factorials [570, 508, 608]. Factors [539, 603, 611]. failure [584, 578, 549]. Failures [16, 381]. Faithful [66]. familial [520]. Fast [145]. fatal [87]. Feature [149]. Fedorov [414]. females [442]. fertilizer [656, 412]. Fibre [161, 360]. Field [386, 482, 546, 438, 96]. fields [542, 107]. financial [648]. Find [386]. Finding [95, 165]. findings [556]. fingers [424]. Fingleton [114]. Finite [56, 564, 604, 6]. First [21]. Fisher [400, 297, 216, 299, 515]. fit [565, 496]. fitted [629, 578, 549]. Fitting [626, 185, 217, 265, 548, 441, 57, 547]. Fixed [346, 95, 543, 345]. flexible [276, 133, 494]. flooding [605]. Flour [162, 187]. fluorescence [602]. fluorescent [521]. Flury [106]. fly [624]. foetal [438]. Folks [45]. food [629, 652]. football [551]. force [424]. forecasting [596]. Forecasts [70]. Forensic [161]. forest [479]. Form [58, 629]. Forms [264]. Formulae [33, 3]. Formulation [413]. Foster [164]. Fourier [145, 532]. Fractals [315]. fractional [508, 394]. fractionally [382]. fractions [616]. frailty [495]. Francis [19]. Frank [389]. freedom [562]. French [48]. Frequencies [514]. Frequency [387, 643, 651]. Front [38, 62, 84, 129, 152, 177, 222, 249, 268, 301, 326, 342, 355, 378, 393, 407, 420, 435, 448, 462, 477, 492, 506, 518, 530]. FRS [260]. Fully [408]. Function [15, 59, 124, 540, 297, 4, 245, 175, 574, 2, 123, 541, 10, 566, 636]. Functional [68, 532, 424]. Functions [185, 29, 352].

G [110, 145, 102, 114, 145]. G.

[317, 68, 114, 52]. G/G/1 [145]. Gajjar [433]. Galambos [49]. Galwey [71]. García [543]. García-Pérez [543]. Gardner [79]. Gauss [460, 440]. Gaussian [45, 542, 344]. General [488, 310, 398, 500, 470, 498, 487, 232, 457, 218, 266, 458, 255, 162, 430, 369, 288, 444, 431, 413, 254, 469, 499, 334, 348, 161, 386]. Generalized [487, 73, 504, 296, 264, 200, 203, 97, 327, 271, 273, 628]. Generating [192, 31]. Generation [472, 98, 514]. genes [585]. genomic [521]. genotype [524]. genotype-environment [524]. GENSTAT [239, 71]. Gentleman [416, 244, 415]. geographical [640]. geophysical [382]. geostatistics [600]. gestational [42]. Geyser [66]. Ghosh [201]. Gibbs [327, 226, 466, 366]. Gifi [234]. GIVE. [55]. Given [149]. GLIM [181, 346, 19]. Gluten [162]. Glynn [164]. Godehart [235]. Goodness [496, 565]. Goodness-of-fit [496, 565]. graph [639]. Graphical [486, 184, 550, 283, 281, 588]. Graphics [240]. Graphs [235, 508, 652]. Great [497]. Greene [376]. Greenfield [171]. Grossman [72]. group [591, 601]. Grouped [324, 346, 56, 594, 381, 284]. Groups [165]. Groves [50]. growth [438, 411]. Guarantee [164]. Guide [292]. Gulag [452]. Gupta [27].

H [376, 47]. H. [315, 68, 168, 239, 141, 139]. haemolytic [328]. Hald [202]. Half [608, 570]. Half-normal [608, 570]. Hand [199]. Handbook [113, 239]. Hardle [236]. harmonics [651]. Harper [24]. Harris [259]. Hasted [25]. Hastie [203]. hazard [569]. Hazards [155, 327, 276, 43, 63]. Hazewinkel [20]. Healey [204]. health [588]. Hearing [498]. height [657]. helmet [330]. Helsinki [166]. Hendry [55]. Hermite [217]. Heron [329]. heterogeneity [595]. Heterogeneous [307]. heteroscedasticity [132]. hidden [623]. Hierarchical [231, 455, 158]. High [323, 216, 299, 365]. Higher [160]. Highest [538]. Hinde [19]. Hinkley [260]. historical [362]. History [202]. Hodgkin [628]. Hollander [576]. Holt [170]. homogeneous [623]. Honour [260]. Hospital [512]. Hours [51]. Human [548, 507, 585, 410, 532, 590, 571, 630, 424, 133]. Hutchinson [291, 292]. Hybrid [446, 471]. hybridization [521]. Hypergeometric [147]. Hypertensives [510].

I. [164, 72]. Idea [30]. identification [364]. Ignorable [279, 659, 606]. II [50]. Illiteracy [210]. image [408, 602]. Images [315, 465, 637, 561]. imaging [532]. immune [610, 494, 423]. Immunodeficiency [507, 410, 571, 630, 133]. impact [572]. Impairment [498, 630]. importance [609]. Imprecise [293]. improve [225]. Improved [271, 384]. Improvement [487, 348]. Imputation [510, 87]. Imputed [334]. In-Process [101]. Inc. [375, 350, 403, 30]. incidence [410, 333]. incidences [362]. Incomplete [59, 124, 4, 1, 2, 119, 527, 123]. incorporate [425]. incorporating [443]. Independence [502, 436]. Independent [126, 287, 515, 125, 191]. Index [36, 579, 621, 661, 64]. indices [275]. individuals [381]. Industrial [255, 201, 455]. Industry [254, 334, 348, 584]. inefficiencies [551]. Inequalities [22]. infant [306]. infected [571]. infection [630, 133]. infections [411]. Infectious [102]. Inference [250, 428, 327, 586]. inflorescences [635]. Influence [169, 630, 229, 127, 613]. influential [271]. Information [85, 178, 269, 356, 421, 478, 531, 70, 400, 29, 631, 522, 564, 604, 214]. Informative [358, 630, 284]. Initial [265]. initiating [423]. injuries [330]. inmate [308]. Innumeracy [210]. inspection [464]. Integer [575]. Integers [298]. Integral [4, 247, 1]. Integrals [353]. intensities [578, 549]. intention [659]. interactions [524]. Interest [488, 310, 398, 500, 470, 498, 487, 232, 457, 458, 255, 162, 430, 369, 288, 444, 431, 413, 254, 469, 499, 334, 348, 161, 386]. interethnic [497]. Interlaboratory [232]. Intermittently [16]. International [375]. interoccasion [534]. interpolation [382]. Interpretation [310, 382]. Interpreting [524, 33, 3, 21]. Interrelated [458]. intersubject [534]. Interval [86, 79]. Intervals [82, 309, 192, 396]. intervention [455]. Introduction [580, 544, 555, 567, 103, 165, 212, 318, 200, 163, 69]. invasion [624]. Inverse [59, 123, 124, 4, 344, 45, 44, 2, 41, 611]. Inversion [554]. Inversions [431]. Investigating [63]. investigation [452, 328]. Investigations [250, 105]. involved [585]. involving [656]. Ireland [182]. Irredundant [120]. irregular [601]. Isotonic [80, 125, 126, 503]. Issues [112]. ITCF [417]. items [88]. Ivanov [107].

J [260, 316, 46, 199, 165, 208, 312, 23, 404, 200, 203, 239, 242, 170, 163, 47, 21, 79, 114]. J. [198, 73, 320, 258, 204, 169, 210, 207, 205, 49, 29, 45, 72, 104, 47, 19, 166, 50, 261]. Jacobi [338]. Jaffer [164]. James [295]. Jenkins [162]. John [237]. Johnson [167]. Johnston [52]. Joint [422]. Jones [164]. Jong [389].

K. [336, 207, 28, 112]. Kaufman [165]. Keedwell [258]. Keep [509]. Keller [47]. Kendall [205]. kernel [493]. Kidney [137]. Kmietowitz [26]. Knif [166]. Know [75]. known [193]. Kotz [167].

L [167, 24, 212, 207, 45, 213, 50, 238]. L. [165, 312, 240, 213, 163, 20, 50]. Labour [77]. Lai [291, 292]. Land [343]. Lane [71]. Langeheine [389]. large [568, 439, 136, 646]. larvae [624]. latent [618]. Latin [258].

Lattice [472]. Lauwerier [315]. Lavit [206]. Lawrence [207]. leaf [602]. Least [360, 265, 244, 415, 416]. leave [309]. leaving [276]. Lee [316, 108]. length [453]. Lengths [81, 526]. Leonenko [107]. Letter [138, 233, 289, 370, 17, 501, 335, 432, 513, 459]. Letters [608, 525, 100, 256, 552, 609, 311, 399]. Level [632, 508]. Levels [186, 442, 654, 436, 74]. Leverage [127]. Lexa [241]. Liepins [317]. Life [312, 453]. Life-Testing [312]. Lifetime [428]. Likelihood [32, 217, 457, 35, 56, 146, 340, 473, 565, 34, 381, 541, 384]. Likelihoods [181, 484]. LIMDEP [376]. Limits [449]. Line [265, 493, 302, 276]. Linear [487, 73, 526, 188, 31, 200, 373, 413, 97, 254, 53, 340, 130, 508, 383, 327, 592, 271, 643, 532, 593, 441, 562, 536, 133, 160, 273, 396, 366, 628]. linearity [383, 278, 89]. lines [651]. LISP [240]. LISP-STAT [240]. Loan [164]. Loans [25]. local [230]. Locating [585]. Location [528]. Log [498, 539, 31, 352, 643, 441, 160]. Log-Concave [352]. Log-Linear [31, 643, 441, 160]. log-non-linear [160]. Log-Normal [498, 539]. Logical [172, 172]. Logistic [400, 44, 195, 359, 396]. Logit [385, 634]. LOGXACT [433]. LOGXACT-TURBO [433]. long [362]. long-term [362]. Longitudinal [358, 571, 630, 601, 606, 645, 641]. Look [570, 66]. loop [278]. Loops [575]. loss [511]. losses [455]. Ltd [262, 405]. Lwin [261]. Lyberg [50]. lymphoma [628]. lynx [593].

M [110, 169, 211, 170, 47, 22, 50, 237, 108, 32]. M. [18, 110, 199, 312, 109, 53, 163, 405, 19, 20]. M/G/1 [110]. MacDonald [336]. Machines [16]. Macros [500, 321]. magnetic [532, 586]. Mahalanobis [485]. making [522]. malaria [616]. males [442]. Malt [369]. Manufacturing [101]. Many [332]. Mapping [134, 532]. Marginal [427, 511]. marginality [609]. Marija [242]. Maritz [261]. Marker [507]. Market [166, 551]. Marking [67]. Markov [626, 617, 547, 623, 534, 365, 598, 507, 572]. Marks [67]. Martens [141]. Martin [79]. Masked [428]. Massey [50]. Matchability [137]. Matched [130, 367, 627]. Matching [137]. Matej [241]. maternal [559, 558, 556, 557]. Mathematica [457]. Mathematical [210, 72, 163, 238]. Mathematics [104]. Matrices [110, 430, 333, 136, 554]. Matrix [400, 6]. Matter [37, 61, 83, 38, 62, 84, 128, 151, 176, 129, 152, 177, 221, 248, 267, 222, 249, 268, 300, 325, 341, 354, 301, 326, 342, 355, 377, 392, 406, 419, 378, 393, 407, 420, 434, 447, 461, 476, 435, 448, 462, 477, 491, 505, 517, 529, 492, 506, 518, 530]. Maurice [205]. Maximum [565, 32, 217, 457, 35, 34, 381, 56, 146, 413, 340, 473]. Maxwell [168]. McCullagh [73]. McRobb [109]. Mead [25]. Means [57, 22, 360, 159]. measured [481]. Measurement [444, 507, 111]. Measurements [150]. Measures [65, 483, 613, 625, 396, 199]. Median [445, 263]. Medical [105, 642]. Mehta [433]. Mendenhall [213, 238]. Messy [52]. Method [95, 253, 150, 8, 93]. Methodology [162, 45, 348, 50]. Methods [251, 111, 357, 305, 290, 312, 430, 105, 51, 237, 617, 41, 647, 582, 547, 571, 642, 497, 261, 68, 54]. Metric [574]. Metropolis [466, 577]. Michael [283]. MICLOG [336]. microscopy [602]. microwave [636]. might [275]. Miguel [543]. Mill [162]. Miller [208]. Milliken [52]. Milling [348]. Million [74]. MIM [349]. Minimization [574]. minimum [309, 538]. mining [593]. Minitab [321, 500, 288]. Miscellanea [44]. Misclassification [280, 521]. misclassified [91]. Mises [274]. Misplaced [449]. Missing [402, 196, 277, 484, 631, 606]. mite [598]. Mitrinovic [22]. Mixed [188, 611].

Mixing [35, 34]. Mixture [56, 281, 467, 597, 656, 584, 495, 365, 564, 604, 616]. Mixtures [344, 345, 565, 198]. ML3 [142]. Model [15, 32, 600, 458, 512, 148, 188, 287, 303, 413, 351, 340, 179, 166, 189, 426, 230, 623, 132, 584, 495, 541, 63, 365, 6, 644, 396, 569, 330, 411, 443]. Model-based [600]. Modelling [319, 487, 363, 260, 344, 612, 551, 255, 638, 16, 367, 314, 308, 11, 481, 460, 440, 276, 413, 254, 331, 635, 19, 65, 161, 427, 617, 154, 624, 640, 365, 394, 588, 662]. Modelling-Non-Linear [487]. Models [73, 279, 90, 220, 235, 56, 312, 31, 209, 444, 97, 127, 385, 99, 130, 409, 467, 626, 465, 484, 327, 524, 597, 547, 618, 511, 271, 131, 13, 643, 648, 593, 441, 578, 549, 590, 496, 425, 158, 276, 43, 40, 535, 598, 133, 642, 160, 273, 564, 604, 569, 366, 628, 70, 113, 200, 203, 106]. modified [384]. modulo [554]. Mogyorodi [72]. Moisture [398]. Mokken [47]. moment [380]. moments [9]. monitoring [588]. monotone [606]. Monro [563, 82, 192]. Monte [617, 192, 547, 534, 497, 572]. Montgomerie [318]. monthly [156]. mortality [11, 643, 134, 306, 91]. motoneurone [586]. motor [578, 549]. motorcycle [330]. movements [598]. moving [541, 594]. Multi [543]. Multi-sets [543]. multidimensional [158]. Multilevel [640]. MULTILOG [375]. Multinomial [514, 409]. Multiple [117, 118, 87, 510, 503, 385, 468, 364]. multiple-changepoint [364]. Multiple-Correlation [117, 118]. Multiple-Imputation [510]. Multiplicative [303, 578, 549]. Multiply [334]. Multitaper [651]. Multivariate [310, 500, 357, 353, 323, 539, 234, 264, 339, 119, 141, 106, 520, 228, 383, 655, 187, 159, 485]. Muraya [295]. Muscatine [597].

N [197, 107]. N. [71, 167, 260, 76, 107, 102]. Naes [141]. Nakagawa [139].

NANOSTAT [322]. Nature [252]. Navigator [30]. Negative [194]. Nelder [73]. Nelson [209, 104]. nematodes [624]. Neolithic [599, 619]. Nested [575]. Nets [169]. networks [596]. neural [596]. neurological [630]. Neuts [110]. Newton [460, 440]. Nicholls [50]. Nigam [27]. Nitin [433]. noise [382]. nominal [273]. Non [487, 540, 279, 245, 632, 60, 280, 475, 532, 373, 431, 413, 53, 52, 592, 623, 648, 593, 536, 283, 160, 659, 278, 566, 89, 606, 366, 284]. Non-Central [245, 60, 475, 540, 566]. non-centrality [566]. Non-Differential [280]. non-homogeneous [623]. Non-Ignorable [279, 659, 606]. non-informative [284]. Non-Linear [373, 413, 53, 532, 592, 593, 536, 366]. non-linearity [278, 89]. non-monotone [606]. non-normality [283]. Non-Replicated [52]. Non-Response [279, 659]. Non-Stationarity [632]. Non-Stationary [53]. non-symmetric [648]. Non-uniqueness [431]. Noncentral [5]. Nonlinear [113, 234, 68]. Nonparametric [446, 607, 486, 157, 236, 593, 54, 159, 423, 561, 615, 10]. Norm [265]. Normal [498, 246, 353, 400, 58, 247, 56, 44, 339, 340, 608, 539, 225, 570]. Normality [474, 383, 283]. Normand [390]. norms [94]. Northern [182]. Norussis [242]. Note [227, 278]. NSURV [374]. nuclear [584]. Null [338]. number [328]. Numbers [543].

O [68]. obesity [597]. Object [240]. Object-Oriented [240]. observation [308]. observational [627]. Observations [361, 271]. observed [193, 89]. obtained [439]. Occupation [334]. occurrence [623]. odds [426, 40]. Odeh [74]. off [302]. off-line [302]. offender [591]. officers [591]. O’Hagan [111]. Oil [386]. Oil-Field [386]. Old [66]. Oliver [169]. One [460, 440, 275]. One-Step [460, 440]. onset [411]. open [278]. Optimal [488, 149, 218, 266, 445, 414].

Optimization [471, 583, 412]. Optimum [442, 182]. Ord [205]. Order [216, 299, 365, 160]. Ordered [279, 401, 514, 131, 160, 330]. ordering [302]. Orderings [80]. Ordinal [223, 607, 12, 496, 40]. Oriented [240]. Orthogonal [446]. Ott [213]. outcome [653]. Outcomes [196, 588]. outlier [228, 485]. Outliers [304, 654]. output [583]. outs [571]. ovary [520]. Overdispersion [227, 568, 590, 658]. Overspecified [458]. Owen [74]. Oxygen [413]. ozone [484].

P [320, 291, 292, 50]. P. [71, 73, 27, 70, 77, 165, 293, 321, 54, 22, 20, 50, 237, 108]. P5.1 [322]. paediatric [494]. Pain [444]. Paired [223, 324, 180, 646, 612, 662, 13]. pairs [559, 193]. PANMARK [389]. papers [451]. parallel [637]. Parameter [646, 23, 351, 166, 368, 418, 193, 566, 412]. Parameters [146, 473, 592]. Parametric [578, 549, 628, 441, 425, 394, 179]. parsimonious [394]. Partial [40, 271]. partially [89]. Particle [190]. Partition [465]. Parts [74]. Patel [433]. Paternity [188]. pathogen [536]. patients [571]. pattern [347, 597, 637]. Patterns [121, 426, 331, 365]. Paul [79]. Paulos [210]. PC [55, 242]. PC-GIVE. [55]. Peace [28, 112]. Peladeau [390]. Penguin [104]. Percentage [95, 122]. Percentile [489]. Pérez [543]. performance [508]. period [308]. Periodograms [157]. Periodontal [252]. Permutation [324]. Permutations [543]. PEST [337]. PEST2.1 [337]. Peter [295]. pharmacokinetic [534]. Pharmacologically [510]. phase [653, 589]. Physics [20]. pill [509]. pilot [438]. pilot-whales [438]. pinch [424]. placebo [560]. placebo-controlled [560]. Plans [209, 479]. plant [547, 536]. plate [536]. plot [634]. Plots [219, 155, 570, 608, 271, 656]. PLR [25].

Plus [143]. Point [181, 305, 95, 347, 9]. Points [472, 489, 95, 282, 122]. Poisson [11, 98, 91]. Pol [389]. Policy [512]. pollution [643]. Polynomial [446, 338, 217]. Polynomials [185, 394]. Population [385, 329, 493, 306, 88, 366]. Population-Averaged [385]. Populations [319, 173]. position [636]. Positive [575]. positron [611]. Possible [490, 174]. Possibly [196, 7]. posterior [538]. Poverty [516, 429]. Power [297]. Practice [316, 290]. Pragmatics [317]. precipitation [623]. Predict [115]. Prediction [398, 135, 656]. predictions [479]. Preference [573]. Prequential [14]. Presence [285, 603]. present [12]. Presentation [130]. prevalence [617]. price [187]. Prices [14]. Priestley [53]. Primary [510]. Principal [470, 282, 136, 106, 613]. Printing [121]. prison [308]. Probabilidades [320]. Probabilities [147, 60, 247, 475, 5, 339, 173, 186, 6, 521, 293]. Probability [111, 320, 353, 212, 202, 49, 44, 72, 163, 605, 276]. probation [591]. probit [330]. Problem [449, 227, 360, 93]. Problemas [320]. Problems [320, 42, 231, 611, 224]. Procedure [86, 265, 410, 563]. Procedures [166, 224]. Process [181, 487, 82, 101, 92, 624, 192, 275, 6]. Processes [539, 29, 197, 20, 629, 441, 191, 313]. Processing [47]. Product [353, 527]. products [153]. profile [384]. Profiles [150, 385]. Profiling [250]. prognosis [653]. Program [489]. Programming [487]. Progression [507, 252]. Project [337]. Projection [463, 610]. propagules [536]. Proportional [155, 43, 327, 40, 63, 569]. proportions [193]. Pseudofactors [225]. Public [334]. Purchasing [109]. Pure [369]. Puri [27]. purpose [556]. Pursuit [463]. Pyzdek [75].

Q [169]. QR [490]. Quadratic [58, 264]. Quality [86, 487, 74, 317, 318, 75, 101, 348, 237, 302, 109]. quantal [224]. Quantile [571, 656, 627]. Quantiles [118, 388, 657]. Quantitatifs [206]. quasi [436]. quasi-independence [436].

R [46, 317, 239, 214]. R. [74, 25, 212, 169, 45, 164, 197, 203, 213, 104, 47, 50, 238]. R19 [4]. R67 [149]. R82 [33]. R83 [59, 123, 124]. R84 [60]. R85 [122]. R86 [147]. R87 [148]. R88 [175]. R89 [247]. R90 [265]. R91 [340]. R92 [388]. R93 [473]. R94 [474]. R95 [475]. R96 [504]. radiation [230]. radiocarbon [397, 614]. Rainfall [523, 644]. Random [232, 98, 29, 107, 542, 189, 618, 182, 494]. Randomization [515, 587]. randomness [274]. range [272]. Rank [388, 99, 634, 230, 404]. Ranking [251]. Rankings [180, 612, 662, 13]. Rapid [324]. rate [584, 438]. Rates [398, 509, 134, 306, 91]. Ratio [59, 124, 4, 2, 385, 123, 426, 590, 182, 159]. ratios [367]. Ratkowsky [113]. RC [32]. Re [130]. Re-Expression [130]. Read [167]. Readings [48]. Reallocation [304]. reanalysed [597]. Reasoning [293]. reconviction [591]. Records [469, 537, 654]. recurrent [468]. Recursive [265, 566]. Reduced [99, 230]. region [656, 538]. Regional [70]. Regions [339, 373, 524]. registered [443]. Regression [446, 307, 298, 458, 80, 126, 113, 560, 323, 504, 207, 388, 296, 68, 373, 413, 127, 394, 503, 195, 359, 484, 456, 125, 230, 657, 131, 615, 132, 648, 496, 571, 309, 6, 642, 384, 628, 91, 236, 208, 257]. Regressions [490, 592, 562, 174]. Regressor [243]. Regressors [99]. Reid [260]. reinforced [92]. Rejection [226, 352, 466, 577]. Related [106, 42]. Relations [68]. Relationship [484, 308]. Relative [516, 429, 628]. release [153]. Reliability [290, 312, 582, 379, 23]. Reliable [64, 67, 275]. Remark [147, 59, 124, 149, 4, 265, 175, 60, 148, 247, 504, 388, 475, 33, 474, 340, 473, 123, 122, 149, 4, 265, 175, 247, 504, 388, 475, 33, 474, 340, 473]. renewal [10]. Renshaw [319]. Repeated [199, 150, 196, 65, 467, 481, 365]. Repetition [543]. replicate [546]. Replicated [52, 533]. replicates [483]. replication [347]. Report [270, 422, 308]. reporting [87]. resampling [642]. Research [144, 204, 112]. Residual [504, 296, 271]. Resolution [345]. resonance [532]. resources [135]. Respiration [398]. Respiratory [548]. Response [285, 279, 348, 409, 442, 607, 154, 629, 496, 40, 536, 659, 652]. responses [439, 224]. restoration [408]. Restricted [97, 484]. Result [194]. retrospective [364]. returns [648]. Reviews [198, 391, 144, 319, 71, 315, 294, 167, 313, 18, 103, 111, 417, 73, 27, 115, 70, 110, 320, 143, 322, 260, 291, 292, 316, 74, 25, 375, 46, 295, 215, 77, 262, 337, 199, 259, 165, 26, 258, 24, 172, 55, 212, 113, 204, 169, 206, 241, 116, 235, 202, 210, 317, 314, 293, 336, 211, 389, 234, 171, 236, 48, 207, 208, 205, 318, 321, 290, 312, 142, 49, 68, 29, 23, 28, 45, 76, 404, 164, 109, 197, 209, 72, 200, 107, 374, 102, 203, 54, 168, 239, 53, 257, 240, 242, 350, 214, 75, 213, 403]. Reviews [376, 170, 30, 104, 163, 105, 47, 101, 349, 69, 405, 19, 112, 433, 141, 78, 21, 140, 106, 79, 22, 51, 114, 52, 20, 166, 50, 139, 237, 108, 390, 261, 238, 201]. rice [613]. Ridge [195]. Ridyard [164]. Rights [450]. Risk [510, 634, 367, 615, 569]. risks [276]. Robbins [82, 192, 563]. Robbins-Monro [82, 192]. Robbins-Monro-based [563]. Robert [294]. Roberts [451]. Robins [232]. Robust [487, 307, 68, 224, 397, 480, 207]. Rodent [303]. Roger [295]. Rolf [389]. Ross [211]. rotation [651]. Round [232]. Rounding [298]. Rousseeuw [165]. Routines [416, 244, 415]. Row [482, 546, 96]. Row-Column [482, 96].

Rows [57]. rule [9]. Rules [454, 609]. Run [81, 526].

S [45, 105, 22]. S.

[167, 25, 211, 48, 164, 168, 140, 20, 201].

sacrifice [535]. Sample

[614, 117, 118, 119, 371, 182]. sampled [284]. sampler [366]. Samples [515, 464, 438]. Sampling [86, 343, 226, 286, 429, 516, 213, 352, 493, 327, 466, 577, 182, 562, 443]. Savas [171]. Saxe [294]. scale [591, 568, 11, 535]. Scaling [581, 298]. Scheaffer [212, 213, 238]. Scheme [183, 164]. Schemes [81]. Science [161, 581]. Sciences [167]. Scientific [375]. Score [86, 183, 591, 284]. Scores [388, 634, 383, 551]. Scoring [305]. Scotch [369]. screening [589]. Sea [632, 545, 186, 654]. Sea-Level [632]. Sea-Levels [186, 654]. Search [82]. Second [71]. Section [488, 580, 544, 555, 567, 310, 398, 500, 470, 498, 487, 232, 457, 458, 255, 162, 430, 369, 288, 444, 431, 413, 254, 469, 499, 334, 348, 161, 386]. sectional [438, 88]. Seed [437, 464]. segment [650]. Seismic [386]. Selecting [173, 550]. Selection [488, 148, 208, 614, 333, 480]. self [308, 89]. self-exciting [89]. self-report [308]. Seminar [313]. Semiparametric [657]. Sequences [387, 454, 650, 626, 88]. Sequential [228, 583, 563]. serial [426]. Series [103, 361, 205, 53, 69, 67, 304, 626, 654, 484, 533, 596, 426, 187, 481, 532, 593, 550, 89, 179]. Set [345, 397]. Sets [575, 120, 99, 543]. Several [372, 401]. sex [590]. sexual [631]. Shaped [339]. shapes [561]. Shewhart [86]. Shewhart-Cumulative [86]. short [654]. Should [75]. Shrinkage [305]. sides [132]. Significance [570, 436]. Simple [217, 265, 380, 386]. Simpler [44]. Simplex [472]. simplify [642]. simplifying [629]. Simpson [239]. SIMSTAT [390]. simulated [360]. Simulating [489, 542]. Simulation [575, 211]. Simultaneous [191].

Single [166, 485]. Singular [430, 425]. Sir [260]. situ [521]. Size [149, 190]. skewed [7]. Skinner [170]. Skyrms [24]. Small [515, 456]. Smirnov [371]. Smith [77, 169, 170]. Smoothing [458, 445, 647, 641]. Smoothly [307]. Snell [260, 46, 239]. Social [26, 204, 581, 568, 306]. Software [391, 144, 294, 417, 115, 143, 322, 375, 295, 215, 582, 262, 337, 172, 55, 241, 116, 336, 389, 171, 142, 404, 374, 242, 350, 403, 376, 30, 349, 405, 433, 78, 79, 390, 116]. Soil [398, 522, 536]. solar [651, 230]. SOLO [262]. Some [154, 587, 275, 66, 187, 592, 631]. Source [346, 345]. Sources [51]. SP. [144]. Space [319, 644]. space-time [644]. spaced [481]. Spatial [615, 386, 114, 286, 331, 274]. spatiotemporal [547]. Specific [385, 612, 662]. Spectral [533, 651]. Spectroscopic [430, 613]. Speeds [363]. Sphericity [338]. Splicing [395]. splines [641]. spongiform [559, 558, 556, 557]. Sprent [54]. SPSS [242, 350]. SPSS/PC [242]. Square [117, 118]. Squares [265, 258, 416, 360, 244, 415]. stable [648]. Stage [196, 194, 437, 379]. Standardization [470]. Standardized [173]. standardizing [657]. Star [339]. Star-Shaped [339]. STAT [417, 240]. STAT-ITCF [417]. State [303, 626, 6]. statements [272]. Stationarity [632]. Stationary [53, 542]. Statistic [338, 504, 296, 283]. Statistical [391, 144, 36, 446, 294, 324, 167, 32, 147, 59, 124, 250, 338, 471, 82, 217, 417, 472, 149, 115, 298, 357, 297, 4, 143, 322, 502, 25, 35, 375, 219, 295, 117, 118, 245, 215, 218, 246, 266, 489, 353, 400, 262, 58, 265, 337, 81, 26, 175, 172, 60, 55, 526, 80, 148, 126, 402, 387, 247, 241, 145, 116, 293, 445, 336, 323, 372, 401, 389, 56, 171, 504, 318, 290, 312, 388, 296, 31, 142, 475, 264, 29, 216, 299, 5, 404, 339, 173, 514, 209, 1, 2, 119, 57, 107, 243, 244, 373, 414, 416, 374]. Statistical [54, 527, 146, 263, 371, 33, 120, 121, 240, 242, 350, 351, 515, 403, 3, 376, 528, 474, 30, 340, 105, 349, 122, 503, 405, 19, 112, 174, 473, 433, 490, 78, 79, 573, 51, 586, 161, 352, 139, 237, 390, 201, 42, 452, 591, 92, 509, 12, 260]. Statistician [292, 519]. Statistics [144, 316, 190, 204, 202, 28, 72, 239, 516, 469, 545, 537, 238, 429, 642, 108, 24, 76, 163, 21, 140]. STATS [404]. STATXACT [78]. Stays [512]. Step [460, 440]. Stephen [79]. stimulation [586]. Stochastic [313, 110, 624, 302, 197, 637, 20, 547, 6]. Stock [166]. Stopping [454, 194]. Straight [265]. Strategies [488, 77]. stratified [182]. stratigraphy [622]. strawberry [635]. Streit [20]. strength [360]. Strict [94]. stroke [569]. Structural [357, 235]. Structure [353, 33, 3, 252, 607, 635]. Structured [110, 655]. Student [264]. Studies [26, 196, 409, 653, 612, 662, 93, 627, 362, 417, 259]. Study [487, 253, 413, 510, 348, 161, 559, 558, 556, 596, 568, 557, 593, 588, 589]. studying [534]. Subject [385, 612, 662, 554]. Subject-Specific [385, 612, 662]. Subjective [254]. subjects [468]. Subsequences [387]. Subset [149, 208, 490, 174]. Subsets [173, 243]. such [626]. Sum [81, 526, 309]. Summands [575]. Summaries [386]. Summarizing [397]. SUPERANOVA [391]. Supplement [416, 244, 415]. Surface [348, 652]. surfaces [629]. Survey [213, 47, 50]. surveys [568, 170]. Survival [90, 220, 385, 272, 381, 43, 535, 423, 284, 628]. survival-sacrifice [535]. Survivorship [259]. symmetric [648]. symmetry [576]. syndrome [610, 617, 328, 494, 423]. SYSTAT [403]. System [428, 87]. Systematic [343]. systems [278, 139].

T [50]. T.

[291, 292, 203, 75, 170, 141, 139, 261].

Tables [502, 26, 148, 402, 372, 401, 277, 57, 436, 158, 160]. tablets [509]. Tableux [206].

Taguchi [348, 508, 253, 8]. Tail [44]. target [456]. targets [153]. Task [344]. Teaching [288]. techniques [534, 550, 625]. technological [379]. technology [629, 652]. Telephone [50]. Temperature [398]. templates [637]. temporal [638]. Term [57, 362]. terminating [423]. Terry [444]. Test [297, 209, 119, 120, 515, 522, 228, 576, 474, 480, 284, 274]. Testing [338, 312, 209, 437, 273, 360, 651, 582, 485]. Tests [324, 383, 371, 166, 565, 496, 436, 158, 384, 278, 89, 379, 589]. Their [110, 202]. Theory [260, 316, 258, 317, 290, 312, 29, 45, 72, 153, 49]. think [275]. those [244, 415, 416]. Three [303, 196, 463, 437, 347]. Three-Dimensional [463, 347]. Three-Stage [196, 437]. Three-State [303]. threshold [278, 89]. threshold-type [278]. Tibshirani [203]. Tierney [240]. Time [103, 279, 596, 361, 53, 362, 69, 385, 67, 304, 484, 533, 272, 426, 187, 532, 593, 550, 63, 644, 89, 179, 423, 319]. Time-Ordered [279]. Times [205, 481]. timing [509]. Tissue [137]. Toker [405]. tomography [611]. Tony [171]. Tool [204]. tools [599, 619]. tooth [511]. Total [329, 484]. Toxicity [409]. transect [493]. Transfer [161]. Transform [145, 132]. Transformation [470]. Transformations [253, 8, 154]. Transformed [310, 130]. transfusion [423]. transition [365, 6]. transmission [558]. Transplantation [137]. Treated [510]. treatment [560, 595]. trees [331]. Trend [180, 526, 362]. trial [560]. Trials [553, 287, 482, 587, 189, 546, 595]. triangular [436]. Trigamma [175]. True [507]. Truncated [56, 9, 423]. truncation [330]. Tufte [214]. Tumorigenicity [303]. tumour [653, 362]. TURBO [433]. Two [343, 126, 119, 277, 57, 371, 194, 515, 99, 379, 508, 42, 653, 647, 125, 546, 603, 625, 616, 589]. two-component [616]. Two-Dimensional [343]. two-level [508]. two-phase [653, 589]. two-replicate [546]. Two-Sample [119, 371]. Two-Stage [194, 379]. Two-Way [277, 57]. Type [398, 110, 444, 278, 89, 274].

U [144, 316]. U-SP. [144]. U-Statistics [316]. UK [254]. Umbrella [80]. Uncertain [188]. Uncertainty [479, 443]. unchanged [309]. unconditional [410]. unemployed [11]. Unemployment [51, 276]. unequal [592]. unequally [481]. unions [497]. uniqueness [431]. UNISTAT [405]. Unit [144]. university [612, 662]. unobservable [581]. unpaired [560]. Unreplicated [570, 608]. Unsupervised [633]. Uppuluri [317]. Uptake [413]. Upton [114]. uraemic [328]. Use [343, 77, 512, 334, 384, 482, 154, 591, 13, 546, 658, 225, 642, 330]. Used [16, 497]. Useful [247]. Using [250, 656, 627, 490, 386, 166, 493, 582, 597, 596, 648, 441, 590, 601, 134, 485, 283, 394, 644, 494, 174, 480, 641, 366, 284, 636, 628]. utilizations [135].

V [260, 107]. V. [27, 317]. vaccine

[647, 495]. Value [346, 400, 425]. Values [74, 265, 504, 296, 277, 156, 485, 606]. Variability [166, 410, 11, 534, 635]. Variable [86, 574, 154, 271, 639]. Variables [279, 332, 458, 58, 264, 243, 351, 480, 550, 625, 40]. Variance [307, 655, 656, 380, 286, 384, 652]. variances [592]. Variate [98]. Variates [126, 125]. Variation [285, 255, 460, 440, 456, 230, 615]. Varietal [488]. Varying [575]. Vasic [22]. vector [541]. Venezuelan [644]. Version [391, 144, 294, 417, 115, 143, 322, 295, 215, 262, 337, 172, 55, 241, 336, 389, 404, 374, 242, 350, 403, 376, 349, 405, 433, 78, 390]. very [42]. Via [145, 583, 327, 179, 382]. View [395]. Vilaplana [320]. Vincze [72]. Virus [507, 410, 571, 630, 133]. Visek [29]. visual [464]. Vol [579, 29, 51, 114, 52]. Volume [85, 178, 269, 356, 421, 478, 531, 620, 660, 661, 621]. voting [659]. Vrancken [47]. Vries [47].

W [26, 321, 47, 237]. W. [24, 236, 209, 72, 213, 47, 101, 50, 238]. Wackerley [238]. Wadley [227]. Wages [51]. Waksberg [50]. Wald [480]. Walley [293]. waning [647]. Water [254]. Watson [504, 296]. Way [277, 57]. Waymire [197]. Ways [96]. Weak [257]. weekly [156]. weight [657]. weighted [594]. weighting [9]. Weiss [76]. Wertz [72]. whales [438]. Wheeler [25]. which [272, 309]. Whiskies [369]. white [382]. Whitten [23]. Wil [389]. Wilks [228]. William [376]. Wilms [653]. Wind [363]. Winter [79]. Wit [47]. within [466]. Wood [77]. Word [15]. Work [51]. Works [30]. Wrongs [450].

Y. [26]. Yannoulis [26]. Yogesh [433].

Z. [26]. Zehna [321].

References

Majumder:1973:SAAa

[1] K. L. Majumder and G. P. Bhattacharjee. Statistical algorithms: Algorithm AS 63: The incomplete beta integral. Applied Statistics, 22(3):409–411, September 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/63. See remark [4].

Majumder:1973:SAAb

[2] K. L. Majumder and G. P. Bhattacharjee. Statistical algorithms: Algorithm AS 64: Inverse of the incomplete beta function ratio. Applied Statistics, 22(3): 411–414, September 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/64. See remark

[4].

Rogers:1973:SAA

[3] C. E. Rogers. Statistical algorithms: Algorithm AS 65: Interpreting structure formulae. Applied Statistics, 22(3): 414–424, September 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/65. See remark

[33].

Cran:1977:SAR

[4] G. W. Cran, K. J. Martin, and G. E. Thomas. Statistical algorithms: Remark AS R19 and Algorithm AS 109: A remark on algorithms: AS 63: The incomplete beta integral AS 64: Inverse of the incomplete beta function ratio. Applied Statistics, 26(1):111–114, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [1, 2, 59].

Lenth:1987:SAA

[5] Russell V. Lenth. Statistical algorithms: Algorithm AS 226: Computing noncentral beta probabilities. Applied Statistics, 36(2):241–244, June 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/226. See

remarks [60, 475].

Robertson:1990:MRM

[6] Chris Robertson. A matrix regression model for the transition probabilities in a finite state stochastic process. Applied Statistics, 39(1):1–19, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1990:EDA

[7] A. C. Kimber. Exploratory data analysis for possibly censored data from skewed distributions. Applied Statistics, 39(1):21–30, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Logothetis:1990:BCT

[8] N. Logothetis. Box–Cox transformations and the Taguchi method. Applied Statistics, 39(1):31–48, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Darvell:1990:WRE

[9] B. W. Darvell. A weighting rule and an end point correction for moments of truncated distributions. Applied Statistics, 39(1):49–54, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schneider:1990:CNE

[10] Helmut Schneider, Bin-Shan Lin, and Colm O’Cinneide. Comparison of nonparametric estimators for the renewal function. Applied Statistics, 39(1):55–61, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Keiding:1990:MEM

[11] Niels Keiding, Per Kragh Andersen, and Kirsten Frederiksen. Modelling excess mortality of the unemployed: choice of scale and extra-Poisson variability. Applied Statistics, 39(1):63–74, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jansen:1990:SAO

[12] J. Jansen. On the statistical analysis of ordinal data when extravariation is present. Applied Statistics, 39(1):75–84, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Joe:1990:EUP

[13] Harry Joe. Extended use of paired comparison models, with application to chess rankings. Applied Statistics, 39 (1):85–93, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bessler:1990:PAC

[14] David A. Bessler and John L. Kling. Prequential analysis of cattle prices. Applied Statistics, 39(1):95–106, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1990:MFW

[15] B. J. R. Bailey. A model for function word counts. Applied Statistics, 39 (1):107–114, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Follmann:1990:MFI

[16] Dean A. Follmann. Modelling failures of intermittently used machines. Applied Statistics, 39(1):115–123, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Minkin:1990:LE

[17] Salomon Minkin and David J. Venzon. Letter to the editors. Applied Statistics, 39(1):125–127, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Blackaby:1990:BRD

[18] D. H. Blackaby. Book reviews: The Distribution of Earnings 1973 to 1986, by M. Adams. Applied Statistics, 39(1):129, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sittampalam:1990:BRS

[19] Yoga Sittampalam. Book reviews: Statistical Modelling in GLIM, by M. Aitkin, D. Anderson, B. Francis and J. Hinde. Applied Statistics, 39(1):130, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittle:1990:BRS

[20] P. Whittle. Book reviews: Stochastic Processes in Physics and Engineering, by S. Albeverio, P. Blanchard, M. Hazewinkel and L. Streit. Applied Statistics, 39(1):130–131, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1990:BRI

[21] A. Springall. Book reviews: Interpreting Data — A First Course in Statistics, by A. J. B. Anderson. Applied Statistics, 39(1):131, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thompson:1990:BRM

[22] J. W. Thompson. Book reviews: Means and Inequalities, by P. S. Bullen, D. S. Mitrinovic and P. M. Vasic. Applied Statistics, 39(1):131–132, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leech:1990:BRP

[23] D. J. Leech. Book reviews: Parameter Estimation in Reliability, by A. C. Cohen and B. J. Whitten. Applied Statistics, 39(1):132–133, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

French:1990:BRC

[24] Simon French. Book reviews: Causation in Decision, Belief Change and Statistics, by W. L. Harper and B. Skyrms. Applied Statistics, 39(1):133–134, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1990:BRP

[25] A. Davies and Alan D. Airth. Book reviews: PLR loans: A statistical exploration, by A. Hasted, R. Mead and S. Wheeler. Applied Statistics, 39(1): 134–135, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Francis:1990:BRS

[26] Brian Francis. Book reviews: Statistical Tables for Economic, Business and Social Studies. 2nd edn, by Z. W. Kmietowitz and Y. Yannoulis. Applied Statistics, 39(1):135–136, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:1990:BRC

[27] G. M. Clarke. Book reviews: Characterizations and Analysis of Block Designs, by A. K. Nigam, P. D. Puri and V. K. Gupta. Applied Statistics, 39 (1):136–137, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Legg:1990:BRB

[28] Lorna Legg. Book reviews: Biopharmaceutical Statistics for Drug Development, by K. E. Peace. Applied Statistics, 39(1):137–138, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:1990:BRI

[29] P. M. Lee. Book reviews: Information Theory, Statistical Decision Functions, Random Processes, Vol. B, by J. A. Visek. Applied Statistics, 39(1):138, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ryder:1990:SSR

[30] Ken Ryder, Edward Brent, and Greta M. Ljung. Statistical software reviews: Statistical Navigator, by The Idea Works Inc. Applied Statistics, 39 (1):139–142, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krajewski:1990:SAA

[31] Pawel Krajewski and Idzi Siatkowski. Statistical algorithms: Algorithm AS 252: Generating classes for log-linear models. Applied Statistics, 39(1): 143–151, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/252.

Becker:1990:SAA

[32] Mark P. Becker. Statistical algorithms: Algorithm AS 253: Maximum likelihood estimation of the RC(M) association model. Applied Statistics, 39 (1):152–167, 1990. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/253.

Payne:1990:SAR

[33] R. W. Payne. Statistical algorithms: Remark AS R82: A remark on Algorithm AS 65: Interpreting structure formulae. Applied Statistics, 39 (1):167–175, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/65. See [3].

Dersimonian:1990:CAM

[34] R. Dersimonian. Correction to Algorithm AS 221: Maximum likelihood estimation of a mixing distribution. Applied Statistics, 39(1):176, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/221.

DerSimonian:1990:SAC

[35] R. DerSimonian. Statistical algorithms: Correction to Algorithm AS 221: Maximum likelihood estimation of a mixing distribution. Applied Statistics, 39(1):176, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1990:ISA

[36] Anonymous. Index of statistical algorithms. Applied Statistics, 39(1): 177–187, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1990:BMPa

[37] Anonymous. Back matter. Applied Statistics, 39(1):[5], 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1990:FMPa

[38] Anonymous. Front matter. Applied Statistics, 39(1):[3], 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

deBragancaPereira:1990:AD

[39] Carlos Alberto de Bragança Pereira and Luis Raúl Pericchi. Analysis of diagnosability. Applied Statistics, 39(2): 189–204, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peterson:1990:PPO

[40] Bercedis Peterson and Frank E. Harrell, Jr. Partial proportional odds models for ordinal response variables. Applied Statistics, 39(2):205–217, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chow:1990:DBC

[41] Shein-Chung Chow and Jun Shao. On the difference between the classical and inverse methods of calibration. Applied Statistics, 39(2):219–228, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bland:1990:ABV

[42] J. M. Bland, J. L. Peacock, H. R. Anderson, O. G. Brooke, and M. De Curtis. The adjustment of birthweight for very early gestational ages: two related problems in statistical analysis. Applied Statistics, 39(2):229–239, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Noura:1990:PHC

[43] A. A. Noura and K. L. Q. Read. Proportional hazards changepoint models in survival analysis. Applied Statistics, 39 (2):241–253, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lin:1990:MSL

[44] Jinn Tyan Lin. Miscellanea: A simpler logistic approximation to the normal tail probability and its inverse. Applied Statistics, 39(2):255–257, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leslie:1990:BRI

[45] J. R. Leslie. Book reviews: The Inverse Gaussian Distribution: Theory, Methodology, and Applications, by R. S. Chhikara and J. L. Folks. Applied Statistics, 39(2):259–260, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Diamond:1990:BRA

[46] Ian Diamond. Book reviews: Analysis of Binary Data. 2nd Edn, by D. R. Cox and E. J. Snell. Applied Statistics, 39 (2):260–261, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Saunders:1990:BRA

[47] D. Saunders. Book reviews: Automation in Survey Processing, by J. H. W. Driessen, W. J. Keller, R. J. Mokken, E. H. J. Vrancken, W. F. M. de Vries and J. W. W. A. Wit. Applied Statistics, 39 (2):261–262, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1990:BRR

[48] Alan Kimber. Book reviews: Readings in Decision Analysis, by S. French. Applied Statistics, 39(2):262, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lane:1990:BRA

[49] John A. Lane. Book reviews: Advanced Probability Theory, by J. Galambos. Applied Statistics, 39(2):263–264, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1990:BRT

[50] David Williams. Book reviews: Telephone Survey Methodology, by R. M. Groves, P. P. Biemer, L. E. Lyberg, J. T. Massey, W. L. Nicholls II and J. Waksberg. Applied Statistics, 39(2):264–265, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Trewsdale:1990:BRS

[51] Janet M. Trewsdale. Book reviews: Statistical Sources and Methods: Vol. 4, Employment, Unemployment, Wages and Hours of Work. Applied Statistics, 39(2):265, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Weatherup:1990:BRA

[52] S. T. C. Weatherup. Book reviews: Analysis of Messy Data, Vol. 2, Non-Replicated Experiments, by G. A. Milliken and D. E. Johnston. Applied Statistics, 39(2):266, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pemberton:1990:BRN

[53] J. Pemberton. Book reviews: Non-Linear and Non-Stationary Time Series Analysis, by M. B. Priestley. Applied Statistics, 39(2):266–268, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moore:1990:BRA

[54] B. A. Moore. Book reviews: Applied Nonparametric Statistical Methods, by P. Sprent. Applied Statistics, 39(2):268, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fry:1990:SSR

[55] Tim Fry and David F. Hendry. Statistical software reviews: PC-GIVE. Version 6.01, by David F. Hendry. Applied Statistics, 39(2):269–272, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1990:SAA

[56] P. N. Jones and G. J. McLachlan. Statistical algorithms: Algorithm AS 254: Maximum likelihood estimation from grouped and truncated data with finite normal mixture models. Applied Statistics, 39(2):273–282, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/254.

Marsh:1990:SAA

[57] N. W. A. Marsh. Statistical algorithms: Algorithm AS 255: Fitting of twoway tables by means for rows, columns and cross-term. Applied Statistics, 39 (2):283–294, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/255.

Farebrother:1990:SAA

[58] R. W. Farebrother. Statistical algorithms: Algorithm AS 256: The distribution of a quadratic form in normal variables. Applied Statistics, 39 (2):294–309, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/256.

Berry:1990:SAA

[59] K. J. Berry, P. W. Mielke, Jr., and G. W. Cran. Statistical algorithms: Algorithm AS R83: A remark on Algorithm AS 109: Inverse of the incomplete beta function ratio. Applied Statistics, 39(2):309–310, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/109. See [4] and correction [123].

Frick:1990:SAA

[60] H. Frick. Statistical algorithms: Algorithm AS R84: A remark on Algorithm AS 226: Computing non-central beta probabilities. Applied Statistics, 39 (2):311–312, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/226. See [5, 475].

Anonymous:1990:BMPb

[61] Anonymous. Back matter. Applied Statistics, 39(2):[4], 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1990:FMPb

[62] Anonymous. Front matter. Applied Statistics, 39(2):[4], 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1990:ITD

[63] A. N. Pettitt and I. Bin Daud. Investigating time dependence in Cox’s proportional hazards model. Applied Statistics, 39(3):313–329, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1990:HRY

[64] A. F. Bissell. How reliable is your capability index? Applied Statistics, 39 (3):331–340, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Verbyla:1990:MRM

[65] A. P. Verbyla and B. R. Cullis. Modelling in repeated measures experiments. Applied Statistics, 39(3):341–356, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Azzalini:1990:LSD

[66] A. Azzalini and A. W. Bowman. A look at some data on the Old Faithful Geyser. Applied Statistics, 39(3):357–365, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Trabelsi:1990:BMT

[67] Abdelwahed Trabelsi and Steven C. Hillmer. Bench-marking time series with reliable bench-marks. Applied Statistics, 39(3):367–379, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawson:1990:BRN

[68] Andrew Lawson. Book reviews: Nonlinear Regression, Functional Relations and Robust Methods, by H. Bunke and G. O. Bunke. Applied Statistics, 39(3):381, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shea:1990:BRA

[69] B. L. Shea. Book reviews: The Analysis of Time Series: An Introduction, 4th edn, by C. Chatfield. Applied Statistics, 39(3):382, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cohen:1990:BRA

[70] Geoff Cohen. Book reviews: Advances in Regional Demography: Information, Forecasts, Models, by P. Congdon and P. Batey. Applied Statistics, 39(3):382–384, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arnold:1990:BRG

[71] Gillian Arnold. Book reviews: Genstat 5: A Second Course, by P. Digby, N. Galwey and P. Lane. Applied Statistics, 39(3):384, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McKenna:1990:BRP

[72] T. J. McKenna. Book reviews: Probability Theory and Mathematical Statistics with Applications, by W. Grossman, J. Mogyorodi, I. Vincze and W. Wertz. Applied Statistics, 39(3):385, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cheek:1990:BRG

[73] P. J. Cheek. Book reviews: Generalized Linear Models, 2nd Edn, by P. McCullagh and J. A. Nelder. Applied Statistics, 39(3):385–386, 1990. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Curram:1990:BRP

[74] J. B. Curram. Book reviews: Parts per Million Values for Estimating Quality Levels, by R. E. Odeh and D. B. Owen. Applied Statistics, 39(3):386–387, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rao:1990:BRW

[75] T. Subba Rao. Book reviews: What Every Engineer Should Know about Quality Control, by T. Pyzdek. Applied Statistics, 39(3):387–388, 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:1990:BRE

[76] Wei Liu. Book reviews: Elementary Statistics, by N. A. Weiss. Applied Statistics, 39(3):388–389, 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1990:BRE

[77] A. G. Evans. Book reviews: Employers’ Labour Use Strategies, by D. Wood and P. Smith. Applied Statistics, 39(3):389, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sprent:1990:SSR

[78] Peter Sprent and Cyrus R. Mehta. Statistical software reviews: STATXACT. version 1.0, by Cytel Software Corporation. Applied Statistics, 39(3):391–395, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swan:1990:SSR

[79] A. V. Swan. Statistical software reviews: Confidence Interval Analysis (CIA), by Martin J. Gardner, Stephen B. Gardner and Paul D. Winter. Applied Statistics, 39(3):395–396, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Geng:1990:SAA

[80] Zhi Geng and Ning-Zhong Shi. Statistical algorithms: Algorithm AS 257: Isotonic regression for umbrella orderings. Applied Statistics, 39(3): 397–402, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/257.

Fellner:1990:SAA

[81] William H. Fellner. Statistical algorithms: Algorithm AS 258: Average run lengths for cumulative sum schemes. Applied Statistics, 39(3): 402–412, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/258.

Buckland:1990:SAA

[82] S. T. Buckland and P. H. Garthwaite. Statistical algorithms: Algorithm AS 259: Estimating confidence intervals by the Robbins-Monro search process. Applied Statistics, 39 (3):413–424, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/259.

Anonymous:1990:BMPc

[83] Anonymous. Back matter. Applied Statistics, 39(3):[4], 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1990:FMPc

[84] Anonymous. Front matter. Applied Statistics, 39(3):[4], 1990. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1990:VIP

[85] Anonymous. Volume information. Applied Statistics, 39(3):[4], 1990. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Amin:1991:VSI

[86] Raid W. Amin and Matoteng M. Ncube. Variable sampling interval combined Shewhart-cumulative score quality control procedure. Applied Statistics, 40 (1):1–12, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heitjan:1991:MIF

[87] Daniel F. Heitjan and Roderick J. A. Little. Multiple imputation for the fatal accident reporting system. Applied Statistics, 40(1):13–29, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1991:EPD

[88] Laurel A. Smith and Denis A. Evans. Estimating a population distribution of sequences of k items from cross-sectional data. Applied Statistics, 40(1):31–42, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tong:1991:TSE

[89] Howell Tong and Iris Yeung. On tests for self-exciting threshold autoregressivetype non-linearity in partially observed time series. Applied Statistics, 40(1):43– 62, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gamerman:1991:DBM

[90] Dani Gamerman. Dynamic Bayesian models for survival data. Applied Statistics, 40(1):63–79, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittemore:1991:PRM

[91] A. S. Whittemore and G. Gong. Poisson regression with misclassified counts: application to cervical cancer mortality rates. Applied Statistics, 40(1):81–93, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1991:SAD

[92] Martin Crowder. A statistical approach to a deterioration process in reinforced concrete. Applied Statistics, 40 (1):95–103, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lewis:1991:PCM

[93] P. A. Lewis, P. W. Jones, J. W. Polak, and H. T. Tillotson. The problem of conversion in method comparison studies. Applied Statistics, 40(1):105–112, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Planitz:1991:SDA

[94] M. Planitz and J. Gates. Strict discrete approximation in the L_{1} and L_{∞} norms. Applied Statistics, 40(1):113–122, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farnum:1991:FPM

[95] Nicholas R. Farnum. A fixed point method for finding percentage points. Applied Statistics, 40(1):123–126, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seeger:1991:WDC

[96] Paul Seeger. Ways of devising complicated row-column designs for field crop experiments. Applied Statistics, 40 (1):127–131, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nyquist:1991:REG

[97] Hans Nyquist. Restricted estimation of generalized linear models. Applied Statistics, 40(1):133–141, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1991:PRV

[98] C. D. Kemp and Adrienne W. Kemp. Poisson random variate generation. Applied Statistics, 40(1):143–158, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Velu:1991:RRM

[99] Raja P. Velu. Reduced rank models with two sets of regressors. Applied Statistics, 40(1):159–170, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Day:1991:LE

[100] S. J. Day, J. P. Royston, J. W. Pratt, Paul D. Haines, J.-T. Lin, and R. W. Farebrother. Letters to the editors. Applied Statistics, 40(1):171–176, 1991.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scaife:1991:BRP

[101] Garrett Scaife and Michael Stuart. Book reviews: In-Process Quality Control for Manufacturing, by W. E. Barkman. Applied Statistics, 40(1):177, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mollison:1991:BRA

[102] Denis Mollison. Book reviews: The Analysis of Infectious Disease Data, by N. G. Becker. Applied Statistics, 40 (1):177–178, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Blight:1991:BRA

[103] P. A. Blight. Book reviews: The Analysis of Time Series: An Introduction. 4th edn, by C. Chatfield. Applied Statistics, 40(1):178–179, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ryrie:1991:BRP

[104] S. C. Ryrie. Book reviews: The Penguin Dictionary of Mathematics, by J. Daintith and R. D. Nelson. Applied Statistics, 40(1):179, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandilands:1991:BRS

[105] Michael H. Sandilands. Book reviews:

Statistical Methods for Medical Investigations, by B. S. Everitt. Applied Statistics, 40(1):179–180, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1991:BRC

[106] Tony Springall. Book reviews: Common Principal Components and Related Multivariate Models, by B. Flury. Applied Statistics, 40(1):180–181, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mellor:1991:BRS

[107] Dale Mellor. Book reviews: Statistical Analysis of Random Fields, by A. V. Ivanov and N. N. Leonenko. Applied Statistics, 40(1):181–182, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wooff:1991:BRB

[108] David Wooff. Book reviews: Bayesian Statistics, by P. M. Lee. Applied Statistics, 40(1):182, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

MacArthur:1991:BRP

[109] Ewan W. MacArthur. Book reviews: Purchasing and Quality, by M. McRobb. Applied Statistics, 40(1):182–183, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conolly:1991:BRS

[110] B. W. Conolly. Book reviews: Structured Stochastic Matrices of M/G/1 Type and Their Applications, by M. F. Neuts. Applied Statistics, 40(1):183–184, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burrell:1991:BRP

[111] Quentin Burrell. Book reviews: Probability: Methods and Measurement, by A. O’Hagan. Applied Statistics, 40 (1):184–185, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1991:BRS

[112] M. K. Smith. Book reviews: Statistical Issues in Drug Research and Development, by K. E. Peace. Applied Statistics, 40(1):185–186, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Giddings:1991:BRH

[113] D. R. Giddings. Book reviews: Handbook of Nonlinear Regression Models, by D. A. Ratkowsky. Applied Statistics, 40 (1):186–187, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watkins:1991:BRS

[114] Alan Watkins. Book reviews: Spatial Data Analysis by Example: Vol 2, Categorical and Directional Data, by G. J. G. Upton and B. Fingleton. Applied Statistics, 40(1):187–188, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coe:1991:SSR

[115] Richard Coe. Statistical software reviews: Predict. Version 1.3. Applied Statistics, 40(1):189–192, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1991:SSR

[116] David J. Hand and Maeve Wall. Statistical software reviews: BMDP 5V, by BMDP Statistical Software. Applied Statistics, 40(1):192–194, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ding:1991:SAAa

[117] Cherng G. Ding and Rolf E. Bargmann. Statistical algorithms: Algorithm AS 260: Evaluation of the distribution of the square of the sample multiplecorrelation coefficient. Applied Statistics, 40(1):195–198, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/260.

Ding:1991:SAAb

[118] Cherng G. Ding and Rolf E. Bargmann. Statistical algorithms: Algorithm AS 261: Quantiles of the distribution of the square of the sample multiplecorrelation coefficient. Applied Statistics, 40(1):199–202, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/261.

Makuch:1991:SAA

[119] Robert W. Makuch, Michael Escobar, and Samuel Merrill III. Statistical algorithms: Algorithm AS 262: A twosample test for incomplete multivariate data. Applied Statistics, 40(1): 202–212, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/262.

Payne:1991:SAAa

[120] R. W. Payne. Statistical algorithms: Algorithm AS 263: Construction of irredundant test sets. Applied Statistics, 40 (1):213–229, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/263.

Payne:1991:SAAb

[121] R. W. Payne. Statistical algorithms:

Algorithm AS 264: Printing of bit patterns. Applied Statistics, 40(1): 229–233, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/264.

Shea:1991:SAA

[122] B. L. Shea. Statistical algorithms: Algorithm AS R85: A remark on Algorithm AS 91: The percentage points of the χ^{2} distribution. Applied Statistics, 40(1):233–235, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/91.

Berry:1991:CAR

[123] K. J. Berry, P. W. Mielke, and G. W.

Cran. Correction of Algorithm AS R83 — A remark on Algorithm AS 109: Inverse of the incomplete beta function ratio. Applied Statistics, 40 (1):236, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/109. See [59].

Berry:1991:SAC

[124] K. J. Berry, P. W. Mielke, Jr., and G. W.

Cran. Statistical algorithms: Correction to Algorithm AS R83 — A remark on Algorithm AS 109: Inverse of the incomplete beta function ratio. Applied Statistics, 40(1):236, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/109.

Geng:1991:CAI

[125] Z. Geng and N.-Z. Shi. Correction to Algorithm AS 206: Isotonic regression in two independent variates. Applied Statistics, 40(1):236, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/206.

Geng:1991:SAC

[126] Zhi Geng and Ning-Zhong Shi. Statistical algorithms: Correction to Algorithm AS 206: Isotonic regression in two independent variates. Applied Statistics, 40(1):236, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Puterman:1991:CLI

[127] M. L. Puterman. Corrigendum: Leverage and influence in autocorrelated regression models. Applied Statistics, 40(1):237, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1991:BMPa

[128] Anonymous. Back matter. Applied Statistics, 40(1):[3], 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1991:FMPa

[129] Anonymous. Front matter. Applied Statistics, 40(1):[4], 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Waddington:1991:MRE

[130] D. Waddington. Matched re-expression for presentation of transformed data in linear models. Applied Statistics, 40 (2):239–248, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holtbrugge:1991:CRM

[131] Werner Holtbrügge and Martin Schumacher. A comparison of regression models for the analysis of ordered categorical data. Applied Statistics, 40 (2):249–259, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kettl:1991:AHT

[132] Stena Kettl. Accounting for heteroscedasticity in the transform both sides regression model. Applied Statistics, 40(2):261–268, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosenberg:1991:BFL

[133] Philip S. Rosenberg and Mitchell H.

Gail. Backcalculation of flexible linear models of the human immunodeficiency virus infection curve. Applied Statistics, 40(2):269–282, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marshall:1991:MDM

[134] Roger J. Marshall. Mapping disease and mortality rates using empirical Bayes estimators. Applied Statistics, 40(2): 283–294, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Benjamini:1991:CCB

[135] Yoav Benjamini and Magid Igbaria.

Clustering categories for better prediction of computer resources utilizations. Applied Statistics, 40(2):295–307, 1991.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Narayanaswamy:1991:PCA

[136] C. R. Narayanaswamy and D. Raghavarao. Principal component analysis of large dispersion matrices. Applied Statistics, 40(2):309–316, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilks:1991:TMM

[137] Walter R. Gilks. Tissue matching and matchability in kidney transplantation. Applied Statistics, 40(2):317–336, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:1991:LE

[138] R. D. Baker and J. B. Tilbury. Letter to the editors. Applied Statistics, 40 (2):337–338, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilson:1991:BRS

[139] G. Tunnicliffe Wilson. Book reviews: Statistical Analysis and Control of Dynamic Systems, by H. Akaike and T. Nakagawa. Applied Statistics, 40(2): 339–340, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1991:BRA

[140] A. Springall. Book reviews: Assignments in Applied Statistics, by S. Conrad. Applied Statistics, 40(2):340–341, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Spjotvoll:1991:BRM

[141] Emil Spjotvoll. Book reviews: Multivariate Calibration, by H. Martens and T. Naes. Applied Statistics, 40 (2):341–342, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kreft:1991:SSR

[142] Ita G. G. Kreft, Kyung-Sung Kim, and Bob Prosser. Statistical software reviews: ML3. Applied Statistics, 40 (2):343–347, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crichton:1991:SSR

[143] Nicola J. Crichton and Jonathan Buckley. Statistical software reviews: EPI-LOG plus. version 2.06. Applied Statistics, 40(2):347–351, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Allin:1991:SSR

[144] Paul Allin and U-SP Development Team. Statistical software reviews: U-SP. Version 3.0, by Applied Statistics Research Unit. Applied Statistics, 40 (2):351–354, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Grubel:1991:SAA

[145] R. Grubel. Statistical algorithms: Algorithm AS 265: G/G/1 via fast Fourier transform. Applied Statistics, 40(2): 355–365, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/265.

Narayanan:1991:SAA

[146] A. Narayanan. Statistical algorithms: Algorithm AS 266: Maximum likelihood estimation of the parameters of the Dirichlet distribution. Applied Statistics, 40(2):365–374, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/266.

Berger:1991:SAA

[147] Roger L. Berger. Statistical algorithms: Algorithm AS R86: A remark on Algorithm AS 152: Cumulative hypergeometric probabilities. Applied Statistics, 40(2):374–375, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/152.

Geng:1991:SAA

[148] Zhi Geng and Xiu-Mei Liu. Statistical algorithms: Algorithm AS R87: A remark on Algorithm AS 185: Automatic model selection in contingency tables. Applied Statistics, 40(2):376, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/185.

Cobby:1991:SAC

[149] John M. Cobby. Statistical algorithms: Correction to Remark AS R67 — A remark on AS 199: A branch and bound algorithm for determining the optimal feature subset of given size. Applied Statistics, 40(2):376–377, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kenward:1991:CMC

[150] M. G. Kenward. Corrigendum: A method for comparing profiles of repeated measurements. Applied Statistics, 40(2):379, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1991:BMPb

[151] Anonymous. Back matter. Applied Statistics, 40(2):[3], 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1991:FMPb

[152] Anonymous. Front matter. Applied Statistics, 40(2):[4], 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shao:1991:CRT

[153] Jun Shao and Shein-Chung Chow. Constructing release targets for drug products: a Bayesian decision theory approach. Applied Statistics, 40(3): 381–390, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Buxton:1991:SCU

[154] J. R. Buxton. Some comments on the use of response variable transformations in empirical modelling. Applied Statistics, 40(3):391–400, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Henderson:1991:APU

[155] Robin Henderson and Alvin Milner.

Aalen plots under proportional hazards. Applied Statistics, 40(3):401–409, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cholette:1991:CAW

[156] Pierre A. Cholette and Norma B. Chhab. Converting aggregates of weekly data into monthly values. Applied Statistics, 40(3):411–422, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Diggle:1991:NCC

[157] Peter J. Diggle and Nicholas I. Fisher. Nonparametric comparison of cumulative periodograms. Applied Statistics, 40 (3):423–434, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1991:ECT

[158] W. Meade Morgan and Brent A. Blumenstein. Exact conditional tests for hierarchical models in multidimensional contingency tables. Applied Statistics, 40(3):435–442, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Padmanabhan:1991:NER

[159] A. R. Padmanabhan. Nonparametric estimation of a ratio of multivariate means and applications. Applied Statistics, 40 (3):443–447, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simonoff:1991:HOE

[160] Jeffrey S. Simonoff and Chih-Ling Tsai. Higher order effects in log-linear and lognon-linear models for contingency tables with ordered categories. Applied Statistics, 40(3):449–458, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wakefield:1991:GIS

[161] J. C. Wakefield, A. M. Skene, A. F. M.

Smith, and I. W. Evett. General interest section: The evaluation of fibre transfer evidence in forensic science: A case study in statistical modelling. Applied Statistics, 40(3):461–476, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fearn:1991:GIS

[162] T. Fearn and P. I. Maris. General interest section: An application of Box-Jenkins methodology to the control of gluten addition in a flour mill. Applied Statistics, 40(3):477–484, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Salt:1991:BRI

[163] D. W. Salt. Book reviews: Introduction to Probability and Mathematical Statistics, by L. J. Bain and M. Engelhardt. Applied Statistics, 40(3):485, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mabbett:1991:BRE

[164] Alan Mabbett. Book reviews: An Evaluation of the Loan Guarantee Scheme, by S. Barrett, D. Colenutt, R. Foster, D. Glynn, S. Jaffer, I. Jones and D. Ridyard. Applied Statistics, 40(3):485–486, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fox:1991:BRF

[165] Wendy R. Fox. Book reviews: Finding Groups in Data: An Introduction to Cluster Analysis, by L. Kaufman and P. J. Rousseeuw. Applied Statistics, 40 (3):486–487, 1991. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Wigodsky:1991:BRP

[166] P. A. Wigodsky. Book reviews: Parameter Variability in the Single Factor Market Model: An Empirical Comparison of Tests and Estimation Procedures Using Data from the Helsinki Stock Exchange, by J. Knif. Applied Statistics, 40(3):487, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baxter:1991:BRE

[167] M. J. Baxter. Book reviews: Encyclopedia of Statistical Sciences, by S. Kotz, N. L. Johnson and C. B. Read. Applied Statistics, 40(3):488, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morrey:1991:BRD

[168] G. H. Morrey. Book reviews: Designing Experiments and Analysing Data, by S. E. Maxwell and H. D. Delaney. Applied Statistics, 40(3):488, 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goodwin:1991:BRI

[169] Paul Goodwin. Book reviews: Influence Diagrams, Belief Nets and Decision Analysis, by R. M. Oliver and J. Q. Smith. Applied Statistics, 40(3):489, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rutherford:1991:BRA

[170] Gilbert Rutherford. Book reviews: Analysis of Complex Surveys, by C. J. Skinner, D. Holt and T. M. F. Smith. Applied Statistics, 40(3):489–490, 1991.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1991:SSR

[171] Byron Jones and Tony Greenfield. Statistical software reviews: DEX, by Tony Greenfield and Deniz Savas. Applied Statistics, 40(3):491–493, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

French:1991:SSR

[172] Simon French. Statistical software reviews: Logical Decision: Version 1.10, by Logical Decisions. Applied Statistics, 40(3):493–494, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lund:1991:SAA

[173] R. E. Lund. Statistical algorithms: Algorithm AS 267: Probabilities and standardized differences for selecting subsets containing the best populations. Applied Statistics, 40(3):495–502, 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/267.

Smith:1991:SAA

[174] D. M. Smith. Statistical algorithms:

Algorithm AS 268: All possible subset regressions using the QR decomposition. Applied Statistics, 40(3): 502–513, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/268. See correction [490].

Francis:1991:SAR

[175] Brian J. Francis. Statistical algorithms: Remark AS R88: A remark on Algorithm AS 121: The trigamma function. Applied Statistics, 40(3): 514–515, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/121.

Anonymous:1991:BMPc

[176] Anonymous. Back matter. Applied Statistics, 40(3):[4], 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1991:FMPc

[177] Anonymous. Front matter. Applied Statistics, 40(3):[5], 1991. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1991:VIP

[178] Anonymous. Volume information. Applied Statistics, 40(3):[4], 1991. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tsay:1992:MCP

[179] Ruey S. Tsay. Model checking via parametric bootstraps in time series analysis. Applied Statistics, 41(1):1–15, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Critchlow:1992:DTP

[180] D. E. Critchlow and J. S. Verducci. Detecting a trend in paired rankings. Applied Statistics, 41(1):17–29, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berman:1992:APP

[181] Mark Berman and T. Rolf Turner. Approximating point process likelihoods with GLIM. Applied Statistics, 41(1):31– 38, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McCallion:1992:OAS

[182] T. McCallion. Optimum allocation in stratified random sampling with ratio estimation as applied to the Northern Ireland December agricultural sample. Applied Statistics, 41(1):39–45, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xiao:1992:CSC

[183] Hongjian Xiao. A cumulative score control scheme. Applied Statistics, 41 (1):47–54, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1992:GAD

[184] David Brown. A graphical analysis of deviance. Applied Statistics, 41(1):55– 62, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Buckland:1992:FDF

[185] Stephen T. Buckland. Fitting density functions with polynomials. Applied Statistics, 41(1):63–76, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tawn:1992:EPE

[186] Jonathan A. Tawn. Estimating probabilities of extreme sea-levels. Applied Statistics, 41(1):77–93, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Grubb:1992:MTS

[187] Howard Grubb. A multivariate time series analysis of some flour price data.

Applied Statistics, 41(1):95–107, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Im:1992:MLM

[188] S. Im. Mixed linear model with uncertain paternity. Applied Statistics, 41 (1):109–116, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ezzet:1992:REM

[189] Farkad Ezzet and John Whitehead. A random effects model for binary data from crossover clinical trials. Applied Statistics, 41(1):117–126, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fieller:1992:SPS

[190] N. R. J. Fieller, E. C. Flenley, and W. Olbricht. Statistics of particle size data. Applied Statistics, 41(1):127–146, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wesolowsky:1992:SAC

[191] G. O. Wesolowsky. Simultaneous acceptance control charts for independent processes. Applied Statistics, 41(1):147–158, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garthwaite:1992:GMC

[192] Paul H. Garthwaite and Stephen T. Buckland. Generating Monte Carlo confidence intervals by the Robbins-Monro process. Applied Statistics, 41 (1):159–171, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OQuigley:1992:EBP

[193] John O’Quigley. Estimating the binomial parameter n on the basis of pairs of known and observed proportions. Applied Statistics, 41(1):173–180, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pepe:1992:TSE

[194] Margaret Sullivan Pepe and Garnet L.

Anderson. Two-stage experimental designs: Early stopping with a negative result. Applied Statistics, 41(1):181–190, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

leCessie:1992:REL

[195] S. le Cessie and J. C. van Houwelingen. Ridge estimators in logistic regression. Applied Statistics, 41(1):191–201, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lipsitz:1992:TSE

[196] Stuart R. Lipsitz, Nan M. Laird, and David P. Harrington. A three-stage estimator for studies with repeated and possibly missing binary outcomes. Applied Statistics, 41(1):203–213, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

MacKenzie:1992:BRS

[197] I. G. MacKenzie. Book reviews:

Stochastic Processes with Applications, by R. N. Bhattacharya and E. C. Waymire. Applied Statistics, 41(1):215, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Agha:1992:BRE

[198] M. Agha. Book reviews: Experiments with Mixtures, by J. A. Cornell. Applied Statistics, 41(1):215–216, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farrow:1992:BRA

[199] Malcolm Farrow. Book reviews: Analysis of Repeated Measures, by M. J. Crowder and D. J. Hand. Applied Statistics, 41(1):216–217, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Meech:1992:BRI

[200] Robert Meech. Book reviews: An Introduction to Generalized Linear Models, by A. J. Dobson. Applied Statistics, 41 (1):217–218, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Young:1992:BRS

[201] Karen Young. Book reviews: Statistical Design and Analysis of Industrial Experiments, by S. Ghosh. Applied Statistics, 41(1):218, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1992:BRH

[202] David J. Hand. Book reviews: A history of probability and statistics and their applications before 1750, by A. Hald. Applied Statistics, 41(1):219, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moon:1992:BRG

[203] Graham Moon. Book reviews: Generalized Additive Models, by T. J. Hastie and R. J. Tibshirani. Applied Statistics, 41(1):219–220, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goda:1992:BRS

[204] David Goda. Book reviews: Statistics:

A Tool for Social Research, 2nd edn, by J. F. Healey. Applied Statistics, 41 (1):220–221, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1992:BRT

[205] Alan Kimber. Book reviews: Times Series, 3rd edn, by Maurice Kendall and J. K. Ord. Applied Statistics, 41(1):221, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1992:BRA

[206] J. C. Gower. Book reviews: Analyse Conjointe de Tableux Quantitatifs, by Ch. Lavit. Applied Statistics, 41(1):222, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1992:BRR

[207] Alan Kimber. Book reviews: Robust Regression — Analysis and Applications, by K. D. Lawrence and J. L. Arthur. Applied Statistics, 41(1):222–223, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1992:BRS

[208] Alan Kimber. Book reviews: Subset Selection in Regression, by A. J. Miller. Applied Statistics, 41(1):223, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Macarthur:1992:BRA

[209] Ewan Macarthur. Book reviews: Accelerated Testing: Statistical Models, Test Plans, and Data Analyses, by W. Nelson. Applied Statistics, 41(1):224–225, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1992:BRI

[210] David Hand. Book reviews: Innumeracy: Mathematical Illiteracy and its Consequences, by J. A. Paulos. Applied Statistics, 41(1):225, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hinde:1992:BRC

[211] John Hinde. Book reviews: A course in simulation, by S. M. Ross. Applied Statistics, 41(1):225–226, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gates:1992:BRI

[212] J. Gates. Book reviews: Introduction to Probability and its Applications, by R. L. Scheaffer. Applied Statistics, 41 (1):226–227, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rawlins:1992:BRE

[213] Ian Rawlins. Book reviews: Elementary Survey Sampling, 4th edn, by R. L. Scheaffer, W. Mendenhall and L. Ott. Applied Statistics, 41(1):227, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rangecroft:1992:BRE

[214] Margaret Rangecroft. Book reviews: Envisioning Information, by E. R. Tufte.

Applied Statistics, 41(1):227–228, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Disney:1992:SSR

[215] John Disney and Alan J. Collins. Statistical software reviews: DESIGN-EASE: Version 1.25 and DESIGN-EXPERT: Version 2.05. Applied Statistics, 41 (1):229–231, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:1992:SAA

[216] Yoong-Sin Lee and Ting-Kwong Lin.

Statistical algorithms: Algorithm AS 269: High order Cornish-Fisher expansion. Applied Statistics, 41(1): 233–240, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/269.

Buckland:1992:SAA

[217] S. T. Buckland. Statistical algorithms:

Algorithm AS 270: Maximum likelihood fitting of Hermite and simple polynomial densities. Applied Statistics, 41 (1):241–266, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/270.

Dunn:1992:SAAa

[218] C. L. Dunn. Statistical algorithms: Algorithm AS 271: General optimal combinatoric classification. Applied Statistics, 41(1):266–274, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/271.

Dewey:1992:SAA

[219] Michael E. Dewey. Statistical algorithms: Algorithm AS 272: Box plots. Applied Statistics, 41(1):274– 284, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/272.

Gamerman:1992:CDB

[220] Dani Gamerman. Corrigendum: Dynamic Bayesian models for survival data. Applied Statistics, 41(1):285, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1992:BMPa

[221] Anonymous. Back matter. Applied Statistics, 41(1):[3], 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1992:FMPa

[222] Anonymous. Front matter. Applied Statistics, 41(1):[4], 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Agresti:1992:AOP

[223] Alan Agresti. Analysis of ordinal paired comparison data. Applied Statistics, 41 (2):287–297, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vidmar:1992:RPD

[224] Thomas J. Vidmar, Joseph W. McKean, and Thomas P. Hettmansperger. Robust procedures for drug combination problems with quantal responses. Applied Statistics, 41(2):299–315, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Monod:1992:PNU

[225] H. Monod and R. A. Bailey. Pseudofactors: normal use to improve design and facilitate analysis. Applied Statistics, 41 (2):317–336, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilks:1992:ARS

[226] W. R. Gilks and P. Wild. Adaptive rejection sampling for Gibbs sampling. Applied Statistics, 41(2):337–348, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1992:NWP

[227] B. J. T. Morgan and D. M. Smith.

A note on Wadley’s problem with overdispersion. Applied Statistics, 41 (2):349–354, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Caroni:1992:SAW

[228] C. Caroni and P. Prescott. Sequential application of Wilks’s multivariate outlier test. Applied Statistics, 41 (2):355–364, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pack:1992:ICA

[229] P. Pack and I. T. Jolliffe. Influence in correspondence analysis. Applied Statistics, 41(2):365–380, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glasbey:1992:RRR

[230] C. A. Glasbey. A reduced rank regression model for local variation in solar radiation. Applied Statistics, 41 (2):381–387, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carlin:1992:HBA

[231] Bradley P. Carlin, Alan E. Gelfand, and Adrian F. M. Smith. Hierarchical Bayesian analysis of changepoint problems. Applied Statistics, 41(2):389–405, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1992:GIS

[232] Martin J. Crowder. General interest section: Interlaboratory comparisons: Round robins with random effects. Applied Statistics, 41(2):409–425, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farebrother:1992:LE

[233] R. W. Farebrother. Letter to the editors. Applied Statistics, 41(2):427, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jolliffe:1992:BRN

[234] Ian Jolliffe. Book reviews: Nonlinear Multivariate Analysis, by A. Gifi. Applied Statistics, 41(2):429–430, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1992:BRG

[235] David J. Hand. Book reviews: Graphs as Structural Models. 2nd Edn, by E. Godehart. Applied Statistics, 41(2):430–431, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1992:BRA

[236] M. C. Jones. Book reviews: Applied Nonparametric Regression, by W. Hardle. Applied Statistics, 41(2):431–432, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wolstenholme:1992:BRS

[237] Linda Wolstenholme. Book reviews:

Statistical Methods in Engineering and Quality Assurance, by P. W. M. John. Applied Statistics, 41(2):432–433, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Young:1992:BRM

[238] Karen Young. Book reviews: Mathematical Statistics with Applications, by W. Mendenhall, D. D. Wackerley and R. L. Scheaffer. Applied Statistics, 41(2):433, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Patterson:1992:BRA

[239] H. D. Patterson. Book reviews: Applied Statistics: A Handbook of GENSTAT Analyses, by E. J. Snell and H. R. Simpson. Applied Statistics, 41(2):433–434, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettit:1992:BRL

[240] Lawrence Pettit. Book reviews: LISP-STAT: An Object-Oriented Environment for Statistical Computing and Dynamic Graphics, by L. Tierney. Applied Statistics, 41(2):434–435, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenhalgh:1992:SSR

[241] David Greenhalgh and Matej Lexa. Statistical software reviews: DYMO. version 1.0, by Matej Lexa. Applied Statistics, 41(2):436–438, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pickles:1992:SSR

[242] Andrew Pickles and Mark Battaglia. Statistical software reviews: SPSS/ PC+: Version 4.0, by Marija J. Norussis. Applied Statistics, 41(2):438–442, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Miller:1992:SAAa

[243] Alan J. Miller. Statistical algorithms: Algorithm AS 273: Comparing subsets of regressor variables. Applied Statistics, 41(2):443–457, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/273.

Miller:1992:SAAb

[244] Alan J. Miller. Statistical algorithms: Algorithm AS 274: Least squares routines to supplement those of Gentleman. Applied Statistics, 41(2): 458–478, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/274. See correction [415].

Ding:1992:SAA

[245] Cherng G. Ding. Statistical algorithms: Algorithm AS 275: Computing the noncentral χ^{2} distribution function. Applied Statistics, 41(2):478–482, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/275.

Dunn:1992:SAAb

[246] C. L. Dunn. Statistical algorithms:

Algorithm AS 276: Normal combinatoric classification. Applied Statistics, 41 (2):483–496, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/276.

Goedhart:1992:SAR

[247] P. W. Goedhart and M. J. W. Jansen.

Statistical algorithms: Remark AS R89: A remark on Algorithm AS 76: An integral useful in calculating central t and bivariate normal probabilities. Applied Statistics, 41(2):496–497, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/76.

Anonymous:1992:BMPb

[248] Anonymous. Back matter. Applied Statistics, 41(2):[4], 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1992:FMPb

[249] Anonymous. Front matter. Applied Statistics, 41(2):[5], 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berry:1992:SIC

[250] D. A. Berry, I. W. Evett, and R. Pinchin.

Statistical inference in crime investigations using deoxyribonucleic acid profiling. Applied Statistics, 41(3):499–531, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bacon-Shone:1992:RMC

[251] John Bacon-Shone. Ranking methods for compositional data. Applied Statistics, 41(3):533–537, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sterne:1992:ANP

[252] J. A. C. Sterne, A. Kingman, and H. Loe. Assessing the nature of periodontal disease progression — an application of covariance structure estimation. Applied Statistics, 41(3):539–552, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fearn:1992:BCT

[253] T. Fearn. Box–Cox transformations and the Taguchi method: An alternative analysis of a Taguchi case study. Applied Statistics, 41(3):553–559, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OHagan:1992:GIS

[254] A. O’Hagan, E. B. Glennie, and R. E. Beardsall. General interest section: Subjective modelling and Bayes linear estimation in the UK water industry. Applied Statistics, 41(3):563–577, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Engel:1992:GIS

[255] J. Engel. General interest section: Modelling variation in industrial experiments. Applied Statistics, 41(3): 579–593, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1992:LE

[256] J. A. Nelder, Stena Kettl, and Harold Ascher. Letters to the editors. Applied Statistics, 41(3):595–600, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettit:1992:BRC

[257] Lawrence Pettit. Book reviews: Conditioning Diagnostics: Collinearity and Weak Data in Regression, by D. A. Belsley. Applied Statistics, 41(3):601, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Franklin:1992:BRL

[258] M. F. Franklin. Book reviews: Latin Squares: New Developments in the Theory and Applications, by J. Denes and A. D. Keedwell. Applied Statistics, 41 (3):601–602, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fayers:1992:BRS

[259] Peter Fayers. Book reviews: Survivorship Analysis for Clinical Studies, by E. K. Harris and A. Albert. Applied Statistics, 41(3):602–603, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1992:BRS

[260] Martin Crowder. Book reviews: Statistical Theory and Modelling: In Honour of Sir David Cox, FRS, by D. V. Hinkley, N. Reid and E. J. Snell. Applied Statistics, 41(3):603–604, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Young:1992:BRE

[261] Karen Young. Book reviews: Empirical Bayes Methods, by J. Maritz and T. Lwin. Applied Statistics, 41(3):604, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1992:SSR

[262] Stephen Evans and Martin Hurley. Statistical software reviews: SOLO: Version 3.0, by BMDP Statistical Software Ltd. Applied Statistics, 41(3):605–609, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Niinimaa:1992:SAA

[263] A. Niinimaa, H. Oja, and J. Nyblom. Statistical algorithms: Algorithm AS 277: The oja bivariate median. Applied Statistics, 41(3):611–617, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/277.

Lecoutre:1992:SAA

[264] Bruno Lecoutre, Jean-Luc Guigues, and Jacques Poitevineau. Statistical algorithms: Algorithm AS 278: Distribution of quadratic forms of multivariate generalized Student variables. Applied Statistics, 41(3):617–627, 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/278.

Farebrother:1992:SAR

[265] R. W. Farebrother. Statistical algorithms: Remark AS R90: Least squares initial values for the L_{1}-norm fitting of a straight line — A remark on Algorithm AS 238: A simple recursive procedure for the L_{1} norm fitting of a straight line. Applied Statistics, 41 (3):627–633, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/238.

Dunn:1992:SAC

[266] C. L. Dunn. Statistical algorithms: Correction to Algorithm AS 271: General optimal combinatoric classification. Applied Statistics, 41(3):634, 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1992:BMPc

[267] Anonymous. Back matter. Applied Statistics, 41(3):[4], 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1992:FMPc

[268] Anonymous. Front matter. Applied Statistics, 41(3):[5], 1992. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1992:VIP

[269] Anonymous. Volume information. Applied Statistics, 41(3):[4], 1992. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1993:RE

[270] D. J. Hand and W. J. Krzanowski. Report of the editors. Applied Statistics, 42(1):1, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hines:1993:IAV

[271] R. J. O’Hara Hines and E. M. Carter.

Improved added variable and partial residual plots for the detection of influential observations in generalized linear models. Applied Statistics, 42(1):3–20, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinse:1993:CSA

[272] Gregg E. Dinse, Walter W. Piegorsch, and Dennis D. Boos. Confidence statements about the time range over which survival curves differ. Applied Statistics, 42(1):21–30, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1993:TAD

[273] Philip J. Smith and Daniel F. Heitjan. Testing and adjusting for departures from nominal dispersion in generalized linear models. Applied Statistics, 42(1):31–41, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zimmerman:1993:BCM

[274] Dale L. Zimmerman. A bivariate Cramer-von Mises type of test for spatial randomness. Applied Statistics, 42 (1):43–54, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kotz:1993:SPC

[275] Samuel Kotz, Wen Lea Pearn, and N. L. Johnson. Some process capability indices are more reliable than one might think. Applied Statistics, 42(1):55–62, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Narendranathan:1993:MPL

[276] Wiji Narendranathan and Mark B. Stewart. Modelling the probability of leaving unemployment: competing risks models with flexible base-line hazards. Applied Statistics, 42(1):63–83, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mandel:1993:ATW

[277] John Mandel. The analysis of twoway tables with missing values. Applied Statistics, 42(1):85–93, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sorour:1993:NTT

[278] A. Sorour and H. Tong. A note on tests for threshold-type non-linearity in open loop systems. Applied Statistics, 42(1):95–104, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conaway:1993:NIN

[279] Mark R. Conaway. Non-ignorable nonresponse models for time-ordered categorical variables. Applied Statistics, 42 (1):105–115, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenland:1993:CND

[280] Sander Greenland and Hermann Brenner. Correcting for non-differential misclassification in ecologic analyses. Applied Statistics, 42(1):117–126, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vining:1993:GAE

[281] G. Geoffrey Vining, John A. Cornell, and Raymond H. Myers. A graphical approach for evaluating mixture designs. Applied Statistics, 42(1):127–138, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Flury:1993:EPP

[282] Bernard D. Flury. Estimation of principal points. Applied Statistics, 42 (1):139–151, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Royston:1993:GDN

[283] Patrick Royston. Graphical detection of non-normality by using Michael’s statistic. Applied Statistics, 42(1):153–158, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wax:1993:STN

[284] Yohanan Wax, Stuart G. Baker, and Blossom H. Patterson. A score test for non-informative censoring using doubly sampled grouped survival data. Applied Statistics, 42(1):159–172, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Boos:1993:ADR

[285] Dennis D. Boos. Analysis of doseresponse data in the presence of extrabinomial variation. Applied Statistics, 42 (1):173–183, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1993:SDE

[286] A. N. Pettitt and A. B. McBratney. Sampling designs for estimating spatial variance components. Applied Statistics, 42(1):185–209, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:1993:ITM

[287] David A. Jackson. Independent trials are a model for disaster. Applied Statistics, 42(1):211–220, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marriott:1993:GIS

[288] John M. Marriott and John C. Naylor. General interest section: Teaching Bayes on MINITAB. Applied Statistics, 42(1):223–232, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fearn:1993:LE

[289] Tom Fearn, Costas Goutis, Jun Shao, and Shein-Chung Chow. Letter to the editors. Applied Statistics, 42(1): 233–234, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1993:BRSa

[290] Alan Kimber. Book reviews: Statistical Methods in Reliability Theory and Practice, by B. D. Bunday. Applied Statistics, 42(1):235–236, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1993:BRC

[291] Martin Crowder. Book reviews: Continuous Bivariate Distributions, Emphasizing Applications, by T. P. Hutchinson and C. D. Lai. Applied Statistics, 42(1):236, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1993:BRE

[292] Martin Crowder. Book reviews: The Engineering Statistician’s Guide to Continuous Bivariate Distributions, by T. P. Hutchinson and C. D. Lai. Applied Statistics, 42(1):237, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1993:BRS

[293] David J. Hand. Book reviews: Statistical Reasoning with Imprecise Probabilities, by P. Walley. Applied Statistics, 42 (1):237–238, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baccini:1993:SSR

[294] Alain Baccini and Robert Saxe. Statistical software reviews: CORANA: Version 1.80, by Robert Saxe. Applied Statistics, 42(1):239–241, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dickson:1993:SSR

[295] Janet Dickson and James Roger. Statistical software reviews: DATACHAIN: Version 2.1, by James Roger and Peter Muraya. Applied Statistics, 42(1): 242–246, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kohn:1993:SAA

[296] Robert Kohn, Thomas S. Shively, and Craig F. Ansley. Statistical algorithms: Algorithm AS 279: Computing p-values for the generalized Durbin-Watson statistic and residual autocorrelations in regression. Applied Statistics, 42(1):249–258, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/279. See [504].

Conlon:1993:SAA

[297] Michael Conlon and Ronald G. Thomas. Statistical algorithms: Algorithm AS 280: The power function for Fisher’s exact test. Applied Statistics, 42(1):

258–260, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/280.

Cole:1993:SAA

[298] T. J. Cole. Statistical algorithms: Algorithm AS 281: Scaling and rounding regression coefficients to integers. Applied Statistics, 42(1):261–268, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/281.

Lee:1993:SAC

[299] Yoong-Sin Lee and Ting-Kwong Lin.

Statistical algorithms: Correction to Algorithm AS 269: High order Cornish-Fisher expansion. Applied Statistics, 42 (1):268–269, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/269.

Anonymous:1993:BMPa

[300] Anonymous. Back matter. Applied Statistics, 42(1):[2], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1993:FMPa

[301] Anonymous. Front matter. Applied Statistics, 42(1):[6], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kirmani:1993:SOA

[302] S. N. U. A. Kirmani and Shyamal Das Peddada. Stochastic ordering approach to off-line quality control. Applied Statistics, 42(2):271–281, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindsey:1993:TSM

[303] Jane C. Lindsey and Louise M. Ryan. A three-state multiplicative model for rodent tumorigenicity experiments. Applied Statistics, 42(2):283–300, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wu:1993:ROT

[304] Lilian Shiao Yen Wu, J. R. M. Hosking, and Nalini Ravishanker. Reallocation outliers in time series. Applied Statistics, 42(2):301–313, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Copas:1993:SPS

[305] J. B. Copas. The shrinkage of point scoring methods. Applied Statistics, 42 (2):315–331, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Quine:1993:EIM

[306] M. P. Quine and S. Quine. Estimation of infant mortality rates categorized by social class for an Australian population. Applied Statistics, 42(2):333–338, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cohen:1993:RSH

[307] Michael Cohen, Siddhartha R. Dalal, and John W. Tukey. Robust, smoothly heterogeneous variance regression. Applied Statistics, 42(2):339–353, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hurrell:1993:MRB

[308] Karen T. Hurrell. Modelling the relationship between crime count and observation period in prison inmates’ selfreport data. Applied Statistics, 42(2): 355–367, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Narula:1993:IWL

[309] Subhash C. Narula, Vince A. Sposito, and John F. Wellington. Intervals which leave the minimum sum of absolute errors regression unchanged. Applied Statistics, 42(2):369–378, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arnold:1993:GIS

[310] G. M. Arnold and A. J. Collins. General interest section: Interpretation of transformed axes in multivariate analysis. Applied Statistics, 42(2):381–400, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:1993:LE

[311] D. Roland Thomas, John Mandel, and M. G. Kenward. Letters to the editors. Applied Statistics, 42(2):401–406, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1993:BRSb

[312] Alan Kimber. Book reviews: Statistical Analysis of Reliability and Life-Testing Models: Theory and Methods. 2nd Edn, by L. J. Bain and M. Engelhardt. Applied Statistics, 42(2):407, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bingham:1993:BRS

[313] Nick Bingham. Book reviews: Seminar on Stochastic Processes, by E. Cinlar.

Applied Statistics, 42(2):408, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1993:BRM

[314] David J. Hand. Book reviews: Modelling Binary Data, by D. Collett. Applied Statistics, 42(2):408–409, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aston:1993:BRF

[315] Philip J. Aston. Book reviews: Fractals — Images of Chaos, by H. Lauwerier. Applied Statistics, 42(2):409, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1993:BRU

[316] Martin Crowder. Book reviews: Ustatistics: Theory and practice, by A. J. Lee. Applied Statistics, 42(2):409–410, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1993:BRD

[317] David J. Hand. Book reviews: Data Quality Control: Theory and Pragmatics, by G. E. Liepins and V. R. R. Uppuluri. Applied Statistics, 42(2):410, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1993:BRI

[318] Alan Kimber. Book reviews: Introduction to Statistical Quality Control. 2nd Edn, by D. C. Montgomerie. Applied Statistics, 42(2):411, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Appleton:1993:BRM

[319] David Appleton. Book reviews: Modelling Biological Populations in Space and Time, by E. Renshaw. Applied Statistics, 42(2):411–412, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conolly:1993:BRP

[320] Brian Conolly. Book reviews: Problemas de Calculo de Probabilidades (Problems in Probability Calculus). 3rd Edn, by J. P. Vilaplana. Applied Statistics, 42(2):412–413, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1993:BRM

[321] Alan Kimber. Book reviews: A Minitab Companion with Macros, by P. W. Zehna. Applied Statistics, 42(2):413, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crichton:1993:SSR

[322] Nicola J. Crichton and M. J. R. Healy.

Statistical software reviews: NANOS-TAT: Version P5.1. Applied Statistics, 42(2):415–419, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkins:1993:SAA

[323] Douglas M. Hawkins and Jeffrey S. Simonoff. Statistical algorithms: Algorithm AS 282: High breakdown regression and multivariate estimation. Applied Statistics, 42(2):423– 432, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/282.

Baker:1993:SAA

[324] R. D. Baker and J. B. Tilbury. Statistical algorithms: Algorithm AS 283: Rapid computation of the permutation paired and grouped t-tests. Applied Statistics, 42(2):432–441, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/283.

Anonymous:1993:BMPb

[325] Anonymous. Back matter. Applied Statistics, 42(2):[4], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1993:FMPb

[326] Anonymous. Front matter. Applied Statistics, 42(2):[6], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dellaportas:1993:BIG

[327] P. Dellaportas and A. F. M. Smith. Bayesian inference for generalized linear and proportional hazards models via Gibbs sampling. Applied Statistics, 42 (3):443–459, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Henderson:1993:ICA

[328] R. Henderson and J. N. S. Matthews. An investigation of changepoints in the annual number of cases of haemolytic uraemic syndrome. Applied Statistics, 42 (3):461–471, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:1993:EAT

[329] G. E. Thomas. Estimating annual total heron population counts. Applied Statistics, 42(3):473–486, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Weiss:1993:BOP

[330] Andrew A. Weiss. A bivariate ordered probit model with truncation: helmet use and motorcycle injuries. Applied Statistics, 42(3):487–499, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Preisler:1993:MSP

[331] Haiganoush K. Preisler. Modelling spatial patterns of trees attacked by bark-beetles. Applied Statistics, 42(3): 501–514, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Denham:1993:CMV

[332] M. C. Denham and P. J. Brown. Calibration with many variables. Applied Statistics, 42(3):515–528, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krzanowski:1993:ASC

[333] W. J. Krzanowski. Attribute selection in correspondence analysis of incidence matrices. Applied Statistics, 42(3):529–541, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schenker:1993:GIS

[334] Nathaniel Schenker, Donald J. Treiman, and Lynn Weidman. General interest section: Analyses of public use decennial census data with multiply imputed industry and occupation codes. Applied Statistics, 42(3):545–556, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robinson:1993:LE

[335] G. K. Robinson. Letter to the editors. Applied Statistics, 42(3):557–558, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Have:1993:SSR

[336] Tom Ten Have. Statistical software reviews: MICLOG: Version 4.1, by K. I. MacDonald. Applied Statistics, 42 (3):559–561, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farrington:1993:SSR

[337] C. P. Farrington, John Whitehead, and Hazel Brunier. Statistical software reviews: PEST2.1: Version 2.1, by PEST Project. Applied Statistics, 42 (3):561–564, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Boik:1993:SAA

[338] Robert J. Boik. Statistical algorithms: Algorithm AS 284: Null distribution of a statistic for testing sphericity and additivity: A Jacobi polynomial expansion. Applied Statistics, 42(3): 567–576, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/284.

Lohr:1993:SAA

[339] Sharon L. Lohr. Statistical algorithms: Algorithm AS 285: Multivariate normal probabilities of starshaped regions. Applied Statistics, 42 (3):576–582, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/285.

Ryder:1993:SAR

[340] W. D. J. Ryder. Statistical algorithms:

Remark AS R91: A remark on Algorithm AS 139: Maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 42(3):583–584, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/139.

Anonymous:1993:BMPc

[341] Anonymous. Back matter. Applied Statistics, 42(3):[2], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1993:FMPc

[342] Anonymous. Front matter. Applied Statistics, 42(3):[6], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunn:1993:TDS

[343] R. Dunn and A. R. Harrison. Twodimensional systematic sampling of land use. Applied Statistics, 42(4):585–601, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Desmond:1993:MTC

[344] A. F. Desmond and G. R. Chapman.

Modelling task completion data with inverse Gaussian mixtures. Applied Statistics, 42(4):603–613, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Renner:1993:RCD

[345] Ross M. Renner. The resolution of a compositional data set into mixtures of fixed source compositions. Applied Statistics, 42(4):615–631, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Engel:1993:AGE

[346] J. Engel. On the analysis of grouped extreme value data with GLIM of fixed source. Applied Statistics, 42 (4):633–640, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baddeley:1993:ATD

[347] A. J. Baddeley, R. A. Moyeed, C. V. Howard, and A. Boyde. Analysis of a three-dimensional point pattern with replication. Applied Statistics, 42(4): 641–668, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tuck:1993:GIS

[348] M. G. Tuck, S. M. Lewis, and J. I. L. Cottrell. General interest section: Response surface methodology and taguchi: A quality improvement study from the milling industry. Applied Statistics, 42(4):671–681, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:1993:SSR

[349] Andy Scott and David Edwards. Statistical software reviews: MIM: Version 2.0, by David Edwards. Applied Statistics, 42(4):683–686, 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Provan:1993:SSR

[350] Frances Provan and Major Lester. Statistical software reviews: SPSS: Version 5, by SPSS Inc. Applied Statistics, 42 (4):686–690, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reilly:1993:SAA

[351] Park M. Reilly, H. V. Reilly, and S. E.

Keeler. Statistical algorithms: Algorithm AS 286: Parameter estimation in the error-in-variables model. Applied Statistics, 42(4):693–701, 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/286.

Wild:1993:SAA

[352] P. Wild and W. R. Gilks. Statistical algorithms: Algorithm AS 287: Adaptive rejection sampling from log-concave density functions. Applied Statistics, 42 (4):701–709, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/287.

Dunnett:1993:SAC

[353] Charles W. Dunnett. Statistical algorithms: Correction to Algorithm AS 251: Multivariate normal probability integrals with product correlation structure. Applied Statistics, 42(4):709, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1993:BMPd

[354] Anonymous. Back matter. Applied Statistics, 42(4):[6], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1993:FMPd

[355] Anonymous. Front matter. Applied Statistics, 42(4):[6], 1993. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1993:VIP

[356] Anonymous. Volume information. Applied Statistics, 42(4):[4], 1993. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coles:1994:SMM

[357] Stuart G. Coles and Jonathan A. Tawn. Statistical methods for multivariate extremes: An application to structural design. Applied Statistics, 43(1):1–48, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). With discussion.

Diggle:1994:IDL

[358] P. Diggle and M. G. Kenward. Informative drop-out in longitudinal data analysis. Applied Statistics, 43(1):49–93, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). With discussion.

leCessie:1994:LRC

[359] S. Le Cessie and J. C. Van Houwelingen. Logistic regression for correlated binary data. Applied Statistics, 43(1):95–108, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1994:LSS

[360] Martin Crowder. Least squares with simulated means for a problem in fibre strength testing. Applied Statistics, 43 (1):109–115, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gudmundsson:1994:TSA

[361] Gudmundur Gudmundsson. Time series analysis of catch-at-age observations. Applied Statistics, 43(1):117–126, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seewald:1994:TTH

[362] Wolfgang Seewald. Time trend in historical controls for tumour incidences in long-term animal studies. Applied Statistics, 43(1):127–137, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coles:1994:DME

[363] Stuart G. Coles and David Walshaw.

Directional modelling of extreme wind speeds. Applied Statistics, 43(1): 139–157, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stephens:1994:BRM

[364] D. A. Stephens. Bayesian retrospective multiple-changepoint identification. Applied Statistics, 43(1):159–178, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Raftery:1994:EMR

[365] Adrian Raftery and Simon Tavare. Estimation and modelling repeated patterns in high order Markov chains with the mixture transition distribution model. Applied Statistics, 43(1):179–199, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wakefield:1994:BAL

[366] J. C. Wakefield, A. F. M. Smith, A. Racine-Poon, and A. E. Gelfand.

Bayesian analysis of linear and nonlinear population models by using the Gibbs sampler. Applied Statistics, 43 (1):201–221, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenland:1994:MRR

[367] Sander Greenland. Modelling risk ratios from matched cohort data: an estimating equation approach. Applied Statistics, 43(1):223–232, 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wiper:1994:BEB

[368] M. P. Wiper and L. I. Pettit. Bayesian estimation of the binomial parameter n. Applied Statistics, 43(1):233–236, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lapointe:1994:GIS

[369] François-Joseph Lapointe and Pierre Legendre. General interest section: A classification of pure malt Scotch whiskies. Applied Statistics, 43(1): 237–257, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawson:1994:LE

[370] Andrew B. Lawson and Dale L. Zimmerman. Letter to the editors. Applied Statistics, 43(1):259–261, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nikiforov:1994:SAA

[371] Andrei M. Nikiforov. Statistical algorithms: Algorithm AS 288: Exact Smirnov two-sample tests for arbitrary distributions. Applied Statistics, 43 (1):265–270, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/288.

Hirji:1994:SAAa

[372] Karim F. Hirji and Stein E. Vollset. Statistical algorithms: Algorithm AS 289: Computing exact distributions for several 2 × 2 tables. Applied Statistics, 43 (1):270–274, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/289.

Miller:1994:SAAa

[373] Alan J. Miller. Statistical algorithms:

Algorithm AS 290: Bivariate confidence regions in non-linear regression. Applied Statistics, 43(1):275–284, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/290.

Milligan:1994:SSR

[374] Marion Milligan and Volker W. Rahlfs.

Statistical software reviews: NSURV: Version 1.0. Applied Statistics, 43(1): 285–287, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dessens:1994:SSR

[375] Jos Dessens and Wim Jansen. Statistical software reviews: MULTILOG, by Scientific Software International, Inc. Applied Statistics, 43(1):287–290, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosenqvist:1994:SSR

[376] Gunnor Rosenqvist and William Greene.

Statistical software reviews: LIMDEP:

Version 6.0, by William H. Greene. Applied Statistics, 43(1):290–293, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:BMPa

[377] Anonymous. Back matter. Applied Statistics, 43(1):[3], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:FMPa

[378] Anonymous. Front matter. Applied Statistics, 43(1):[5], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whitmore:1994:TSR

[379] G. A. Whitmore, K. D. S. Young, and A. C. Kimber. Two-stage reliability tests with technological evolution: a Bayesian analysis. Applied Statistics, 43 (2):295–307, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lipsitz:1994:SME

[380] Stuart R. Lipsitz, Nan M. Laird, and Troyen A. Brennan. Simple moment estimates of the κ-coefficient and its variance. Applied Statistics, 43(2):309–323, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huet:1994:MLE

[381] S. Huet and A. Kaddour. Maximum likelihood estimation in survival analysis with grouped data on censored individuals and continuous data on failures. Applied Statistics, 43(2):325–333, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walden:1994:IGB

[382] A. T. Walden. Interpretation of geophysical borehole data via interpolation of fractionally differenced white noise. Applied Statistics, 43(2):335–345, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1994:TLM

[383] D. R. Cox and Nanny Wermuth. Tests of linearity, multivariate normality and the adequacy of linear scores. Applied Statistics, 43(2):347–355, 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simonoff:1994:UMP

[384] Jeffrey S. Simonoff and Chih-Ling Tsai.

Use of modified profile likelihood for improved tests of constancy of variance in regression. Applied Statistics, 43 (2):357–370, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

TenHave:1994:SSP

[385] Thomas R. Ten Have and David H. Uttal. Subject-specific and populationaveraged continuation ratio logit models for multiple discrete time survival profiles. Applied Statistics, 43(2):371–384, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walden:1994:GIS

[386] A. T. Walden. General interest section:

Spatial clustering: Using simple summaries of seismic data to find the edge of an oil-field. Applied Statistics, 43 (2):385–398, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gentleman:1994:SAA

[387] Jane F. Gentleman. Statistical algorithms: Algorithm AS 291: The distribution of the frequency of subsequences in alphabetic sequences, as exemplified by deoxyribonucleic acid. Applied Statistics, 43(2):401–409, 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/291.

Koenker:1994:SAR

[388] Roger Koenker and Vasco d’Orey. Statistical algorithms: Remark AS R92: A remark on Algorithm AS 229: Computing dual regression quantiles and regression rank scores. Applied Statistics, 43 (2):410–414, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/229.

Humphreys:1994:SSR

[389] K. Humphreys, C. J. Skinner, Frank van de Pol, and Rolf Langeheine. Statistical software reviews: PANMARK: Version 2.2.6, by Frank van de Pol, Rolf Langeheine and Wil de Jong. Applied Statistics, 43(2):415–417, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wooff:1994:SSR

[390] David Wooff and Normand Peladeau. Statistical software reviews: SIMSTAT: Version 2.0, by Normand Peladeau. Applied Statistics, 43(2):417–422, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitkin:1994:SSR

[391] Irit Aitkin, Murray Aitkin, and Clifford Baron. Statistical software reviews:

SUPERANOVA: Version 1.11, by ABA-CUS Concepts. Applied Statistics, 43 (2):422–427, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:BMPb

[392] Anonymous. Back matter. Applied Statistics, 43(2):[3], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:FMPb

[393] Anonymous. Front matter. Applied Statistics, 43(2):[5], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Royston:1994:RUF

[394] Patrick Royston and Douglas G. Altman. Regression using fractional polynomials of continuous covariates: parsimonious parametric modelling. Applied Statistics, 43(3):429–467, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1994:AVS

[395] J. A. Nelder. An alternative view of the splicing data. Applied Statistics, 43 (3):469–476, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanEwijk:1994:CMC

[396] P. H. Van Ewijk and J. A. Hoekstra.

Curvature measures and confidence intervals for the linear logistic model. Applied Statistics, 43(3):477–487, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Christen:1994:SSR

[397] J. Andres Christen. Summarizing a set of radiocarbon determinations: a robust approach. Applied Statistics, 43 (3):489–503, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Azzalini:1994:GIS

[398] Adelchi Azzalini and Peter J. Diggle. General interest section: Prediction of soil respiration rates from temperature, moisture content and soil type. Applied Statistics, 43(3):505–526, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vanderHeijden:1994:LE

[399] Peter G. M. van der Heijden, John Mandel, and C. A. Glasbey. Letters to the editors. Applied Statistics, 43(3):527–530, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Escobar:1994:SAA

[400] Luis A. Escobar and William Q. Meeker, Jr. Statistical algorithms: Algorithm AS 292: Fisher information matrix for the extreme value, normal and logistic distributions and censored data. Applied Statistics, 43(3):533–540, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/292.

Hirji:1994:SAAb

[401] Karim F. Hirji and Stein E. Vollset. Statistical algorithms: Algorithm AS 293: Computing exact distributions for several ordered 2×K tables. Applied Statistics, 43(3):541–548, 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/293.

Geng:1994:SAA

[402] Zhi Geng, Chooichiro Asano, Minoru Ichimura, and Hiroshi Kimura. Statistical algorithms: Algorithm AS 294: Decomposability and collapsibility for contingency tables with missing data. Applied Statistics, 43(3): 548–554, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/294.

Robinson:1994:SSR

[403] Rachel Robinson and Ana Timberlake.

Statistical software reviews: SYSTAT: Version 5.02, by SYSTAT Inc. Applied Statistics, 43(3):555–560, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:1994:SSR

[404] Wei Liu and D. J. Best. Statistical software reviews: RANK STATS: Version 1.2, by D. J. Best. Applied Statistics, 43 (3):560–562, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shoesmith:1994:SSR

[405] Eddie Shoesmith and M. A. Toker. Statistical software reviews: UNISTAT: Version 1.2, by M. A. Toker and Unistat Ltd. Applied Statistics, 43(3):562–568, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:BMPc

[406] Anonymous. Back matter. Applied Statistics, 43(3):[4], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:FMPc

[407] Anonymous. Front matter. Applied Statistics, 43(3):[5], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heikkinen:1994:FBA

[408] Juha Heikkinen and Harri Högmander. Fully Bayesian approach to image restoration with an application in biogeography. Applied Statistics, 43(4): 569–582, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhu:1994:DRM

[409] Y. Zhu, D. Krewski, and W. H. Ross. Dose-response models for correlated multinomial data from developmental toxicity studies. Applied Statistics, 43(4):583–598, 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hay:1994:PEU

[410] Joel W. Hay and Frank A. Wolak. A procedure for estimating the unconditional cumulative incidence curve and its variability for the human immunodeficiency virus. Applied Statistics, 43(4):599–624, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittaker:1994:DCM

[411] Joe Whittaker and Sylvia Frühwirth-Schnatter. A dynamic changepoint model for detecting the onset of growth in bacteriological infections. Applied Statistics, 43(4):625–640, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wallach:1994:EPE

[412] D. Wallach and P. Loisel. Effect of parameter estimation on fertilizer optimization. Applied Statistics, 43(4): 641–651, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nevill:1994:GIS

[413] Alan M. Nevill and Roger L. Holder.

General interest section: Modelling maximum oxygen uptake — A casestudy in non-linear regression model formulation and comparison. Applied Statistics, 43(4):653–666, 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Miller:1994:SAAb

[414] Alan J. Miller and Nam-Ky Nguyen.

Statistical algorithms: Algorithm AS 295: A fedorov exchange algorithm for D-optimal design. Applied Statistics, 43 (4):669–677, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/295.

Miller:1994:CAL

[415] A. J. Miller. Correction to Algorithm AS 274: Least squares routines to supplement those of Gentleman. Applied Statistics, 43(4):678, 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/274. See [244].

Miller:1994:SAC

[416] Alan J. Miller. Statistical algorithms:

Correction to Algorithm AS 274: Least squares routines to supplement those of gentleman. Applied Statistics, 43(4):678, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Butler:1994:SSR

[417] Ruth Butler and J. Tranchefort. Statistical software reviews: STAT-ITCF: Version 4, by Department of Statistical Studies. Applied Statistics, 43 (4):679–682, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wiper:1994:CBE

[418] M. P. Wiper and L. I. Pettit. Corrigendum: Bayesian estimation of the binomial parameter n. Applied Statistics, 43(4):685, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:BMPd

[419] Anonymous. Back matter. Applied Statistics, 43(4):[3], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:FMPd

[420] Anonymous. Front matter. Applied Statistics, 43(4):[5], 1994. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1994:VIP

[421] Anonymous. Volume information. Applied Statistics, 43(4):[3], 1994. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krzanowski:1995:RJE

[422] W. J. Krzanowski and D. A. Preece. Report of the joint editors. Applied Statistics, 44(1):1, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tu:1995:NES

[423] Xin Ming Tu. Nonparametric estimation of survival distributions with censored initiating time, and censored and truncated terminating time: application to transfusion data for acquired immune deficiency syndrome. Applied Statistics, 44(1):3–16, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ramsay:1995:FDA

[424] J. O. Ramsay, X. Wang, and R. Flanagan. A functional data analysis of the pinch force of human fingers. Applied Statistics, 44(1):17–30, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Milan:1995:APB

[425] Luis Milan and Joe Whittaker. Application of the parametric bootstrap to models that incorporate a singular value decomposition. Applied Statistics, 44(1):31–49, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fitzmaurice:1995:MBT

[426] Garrett M. Fitzmaurice and Stuart R.

Lipsitz. A model for binary time series data with serial odds ratio patterns. Applied Statistics, 44(1):51–61, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Balagtas:1995:MMC

[427] Cecile C. Balagtas, Mark P. Becker, and Joseph B. Lang. Marginal modelling of categorical data from crossover experiments. Applied Statistics, 44(1):63–77, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reiser:1995:BIM

[428] B. Reiser, I. Guttman, Dennis K. J. Lin, Frank M. Guess, and John S. Usher. Bayesian inference for masked system lifetime data. Applied Statistics, 44 (1):79–90, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Preston:1995:SDR

[429] Ian Preston. Sampling distributions of relative poverty statistics. Applied Statistics, 44(1):91–99, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [516].

Krzanowski:1995:GIS

[430] W. J. Krzanowski, P. Jonathan, W. V. McCarthy, and M. R. Thomas. General interest section: Discriminant analysis with singular covariance matrices: Methods and applications to spectroscopic data. Applied Statistics, 44(1): 101–115, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1995:GIS

[431] Byron J. T. Morgan and Andrew P. G. Ray. General interest section: Nonuniqueness and inversions in cluster analysis. Applied Statistics, 44(1): 117–134, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shirley:1995:LE

[432] Eryl A. C. Shirley. Letter to the editors. Applied Statistics, 44(1):135–137, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1995:SSR

[433] Peter W. F. Smith and Cyrus R.

Mehta. Statistical software reviews: LOGXACT-TURBO: Version 1.1, by Yogesh Gajjar, Cyrus Mehta and Nitin Patel. Applied Statistics, 44(1):139–141, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:BMPa

[434] Anonymous. Back matter. Applied Statistics, 44(1):[3], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:FMPa

[435] Anonymous. Front matter. Applied Statistics, 44(1):[4], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McDonald:1995:ECT

[436] J. W. McDonald and P. W. F.

Smith. Exact conditional tests of quasiindependence for triangular contingency tables: estimating attained significance levels. Applied Statistics, 44(2):143–151, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ridout:1995:TSD

[437] M. S. Ridout. Three-stage designs for seed testing experiments. Applied Statistics, 44(2):153–162, 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rothery:1995:EFG

[438] P. Rothery, A. R. Martin, L. Hiby, and S. Garside. Estimating foetal growth rate from cross-sectional samples of field data with application to pilot-whales. Applied Statistics, 44(2):163–172, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Miller:1995:ACR

[439] Michael E. Miller. Analysing categorical responses obtained from large clusters. Applied Statistics, 44(2):173–186, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morel:1995:OSG

[440] Jorge G. Morel and Kenneth J. Koehler. A one-step Gauss–Newton estimator for modelling categorical data with extraneous variation. Applied Statistics, 44 (2):187–200, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [460].

Lindsey:1995:FPC

[441] J. K. Lindsey. Fitting parametric counting processes by using log-linear models. Applied Statistics, 44(2):201–212, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:1995:ODL

[442] A. C. Atkinson, C. G. B. Demetrio, and S. S. Zocchi. Optimum dose levels when males and females differ in response. Applied Statistics, 44(2):213–226, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

York:1995:BDR

[443] Jeremy York, David Madigan, Ivar Heuch, and Rolv Terje Lie. Birth defects registered by double sampling: a Bayesian approach incorporating covariates and model uncertainty. Applied Statistics, 44(2):227–242, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Matthews:1995:GIS

[444] J. N. S. Matthews and K. P. Morris.

General interest section: An application of Bradley-Terry-type models to the measurement of pain. Applied Statistics, 44(2):243–255, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hardle:1995:SAA

[445] W. Härdle and W. Steiger. Statistical algorithms: Algorithm AS 296: Optimal median smoothing. Applied Statistics, 44(2):258–264, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/296.

Azari:1995:SAA

[446] A. S. Azari and H. G. Müller. Statistical algorithms: Algorithm AS 297: Orthogonal polynomial and hybrid estimators for nonparametric regression. Applied Statistics, 44(2):264–268, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/297.

Anonymous:1995:BMPb

[447] Anonymous. Back matter. Applied Statistics, 44(2):[2], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:FMPb

[448] Anonymous. Front matter. Applied Statistics, 44(2):[5], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Alwan:1995:PMC

[449] Layth C. Alwan and Harry V. Roberts. The problem of misplaced control limits. Applied Statistics, 44(3):269–278, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Caulcutt:1995:RWC

[450] Roland Caulcutt. The rights and wrongs of control charts. Applied Statistics, 44 (3):279–288, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:DPA

[451] Anonymous. Discussion of the papers by Alwan and Roberts, and Caulcutt. Applied Statistics, 44(3):289–306, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Blyth:1995:DGE

[452] Stephen Blyth. The dead of the Gulag: An experiment in statistical investigation. Applied Statistics, 44(3):307–321, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ramon:1995:AEA

[453] Jose Ramon, G. Albert, and Laurence A. Baxter. Applications of the EM algorithm to the analysis of life length data. Applied Statistics, 44(3):323–341, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilmour:1995:SRS

[454] S. G. Gilmour and R. Mead. Stopping rules for sequences of factorial designs. Applied Statistics, 44(3):343–355, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arjas:1995:ALC

[455] Elja Arjas and Liping Liu. Assessing the losses caused by an industrial intervention: A hierarchical Bayesian approach. Applied Statistics, 44(3):357–368, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chan:1995:RAD

[456] L. K. Chan and T. K. Mak. A regression approach for discovering small variation around a target. Applied Statistics, 44 (3):369–377, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Currie:1995:GIS

[457] Iain D. Currie. General interest section: Maximum likelihood estimation and mathematica. Applied Statistics, 44 (3):379–394, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Elston:1995:GIS

[458] D. A. Elston and M. F. Proe. General interest section: Smoothing regression coefficients in an overspecified regression model with interrelated explanatory variables. Applied Statistics, 44 (3):395–406, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilson:1995:LE

[459] S. R. Wilson and J. A. Christen. Letter to the editors. Applied Statistics, 44 (3):407–409, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morel:1995:COS

[460] Jorge G. Morel and Kenneth J. Koehler. Corrigendum: A one-step Gauss– Newton estimator for modelling categorical data with extraneous variation. Applied Statistics, 44(3):410, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [440].

Anonymous:1995:BMPc

[461] Anonymous. Back matter. Applied Statistics, 44(3):[2], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:FMPc

[462] Anonymous. Front matter. Applied Statistics, 44(3):[4], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nason:1995:TDP

[463] Guy Nason. Three-dimensional projection pursuit. Applied Statistics, 44 (4):411–430, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitken:1995:CDV

[464] C. G. G. Aitken, J. Shaw, and M. Talbot. Contaminant detection in the visual inspection of seed samples. Applied Statistics, 44(4):431–440, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aykroyd:1995:PMA

[465] R. G. Aykroyd. Partition models in the analysis of autoradiographic images. Applied Statistics, 44(4):441–454, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilks:1995:ARM

[466] W. R. Gilks, N. G. Best, and K. K. C.

Tan. Adaptive rejection Metropolis sampling within Gibbs sampling. Applied Statistics, 44(4):455–472, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [577].

vanDuijn:1995:MMA

[467] Marijtje A. J. van Duijn and Ulf Bockenholt. Mixture models for the analysis of repeated count data. Applied Statistics, 44(4):473–485, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawless:1995:ARE

[468] J. F. Lawless. The analysis of recurrent events for multiple subjects. Applied Statistics, 44(4):487–498, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robinson:1995:GIS

[469] Michael E. Robinson and Jonathan A.

Tawn. General interest section: Statistics for exceptional athletics records. Applied Statistics, 44(4):499–511, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baxter:1995:GIS

[470] M. J. Baxter. General interest section:

Standardization and transformation in principal component analysis, with applications to archaeometry. Applied Statistics, 44(4):513–527, 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brooks:1995:SAA

[471] Stephen P. Brooks. Statistical algorithms: Algorithm AS 298: A hybrid optimization algorithm. Applied Statistics, 44(4):530–533, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/298.

Chasalow:1995:SAA

[472] Scott D. Chasalow and Richard J. Brand. Statistical algorithms: Algorithm AS 299: Generation of simplex lattice points. Applied Statistics, 44 (4):534–545, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/299.

Smith:1995:SAR

[473] D. M. Smith and M. S. Ridout. Statistical algorithms: Remark AS R93: A remark on Algorithm AS 189: Maximum likelihood estimation of the parameters of the beta binomial distribution. Applied Statistics, 44(4): 545–547, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/189.

Royston:1995:SAR

[474] Patrick Royston. Statistical algorithms: Remark AS R94: A remark on Algorithm AS 181: The W -test for normality. Applied Statistics, 44(4): 547–551, 1995. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/181.

Lam:1995:SAR

[475] Ming Long Lam. Statistical algorithms:

Remark AS R95: A remark on Algorithm AS 226: Computing non-central beta probabilities. Applied Statistics, 44 (4):551–552, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/226. See [5, 60].

Anonymous:1995:BMPd

[476] Anonymous. Back matter. Applied Statistics, 44(4):[2], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:FMPd

[477] Anonymous. Front matter. Applied Statistics, 44(4):[5], 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1995:VIP

[478] Anonymous. Volume information. Applied Statistics, 44(4):[4], 1995. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Alho:1996:UEP

[479] Juha M. Alho, Jyrki Kangas, and Osmo Kolehmainen. Uncertainty in expert predictions of the ecological consequences of forest plans. Applied Statistics, 45(1):1– 14, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sommer:1996:VSU

[480] Suzanne Sommer and Richard M. Huggins. Variables selection using the Wald test and a robust CP. Applied Statistics, 45(1):15–29, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:1996:MRS

[481] Phillipe Lambert. Modelling of repeated series of count data measured at unequally spaced times. Applied Statistics, 45(1):31–38, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1996:RCF

[482] E. R. Williams and J. A. John. Rowcolumn factorial designs for use in agricultural field trials. Applied Statistics, 45(1):39–46, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:1996:DMA

[483] Christine M. Anderson and C. F. Jeff Wu. Dispersion measures and analysis for factorial directional data with replicates. Applied Statistics, 45(1):47–61, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Basu:1996:RBM

[484] Sabyasachi Basu and Gregory C. Reinsel. Relationship between missing data likelihoods and complete data restricted likelihoods for regression time series models: An application to total ozone data. Applied Statistics, 45(1):63–72, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Penny:1996:ACV

[485] Kay I. Penny. Appropriate critical values when testing for a single multivariate outlier by using the Mahalanobis distance. Applied Statistics, 45(1):73–81, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bowman:1996:GCN

[486] Adrian Bowman and Stuart Young.

Graphical comparison of nonparametric curves. Applied Statistics, 45(1):83–98, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brinkley:1996:GIS

[487] P. A. Brinkley, K. P. Meyer, and J. C.

Lu. General interest section: Combined generalized linear modelling-non-linear programming approach to robust process design — A case-study in circuit board quality improvement. Applied Statistics, 45(1):99–110, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Andrews:1996:GIS

[488] N. J. Andrews and R. N. Curnow. General interest section: Optimal strategies for varietal selection. Applied Statistics, 45(1):111–125, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunn:1996:SAA

[489] C. L. Dunn. Statistical algorithms: Algorithm AS 300: An efficient program for simulating percentile points. Applied Statistics, 45(1):129–133, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/300.

Smith:1996:SAC

[490] D. M. Smith. Statistical algorithms:

Correction to Algorithm AS 268: All possible subset regressions using the QR decomposition. Applied Statistics, 45(1):134, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/268. See [174].

Anonymous:1996:BMPa

[491] Anonymous. Back matter. Applied Statistics, 45(1):[2], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1996:FMPa

[492] Anonymous. Front matter. Applied Statistics, 45(1):[5], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:1996:KED

[493] Song Xi Chen. A kernel estimate for the density of a biological population by using line transect sampling. Applied Statistics, 45(2):135–150, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shi:1996:APC

[494] Minggao Shi, Robert E. Weiss, and Jeremy M. G. Taylor. An analysis of paediatric CD4 counts for acquired immune deficiency syndrome using flexible random curves. Applied Statistics, 45 (2):151–163, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Longini:1996:FMM

[495] Ira M. Longini, Jr. and M. Elizabeth Halloran. A frailty mixture model for estimating vaccine efficacy. Applied Statistics, 45(2):165–173, 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lipsitz:1996:GFT

[496] Stuart R. Lipsitz, Garrett M. Fitzmaurice, and Geert Molenberghs. Goodnessof-fit tests for ordinal response regression models. Applied Statistics, 45 (2):175–190, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1996:MCE

[497] Peter W. F. Smith, John W. McDonald, Jonathan J. Forster, and Ann M. Berrington. Monte Carlo exact methods used for analysing interethnic unions in Great Britain. Applied Statistics, 45 (2):191–202, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bowater:1996:GIS

[498] R. J. Bowater, J. B. Copas, and O. A.

Machado. General interest section: Hearing impairment and the log-normal distribution. Applied Statistics, 45(2): 203–217, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sasieni:1996:GIS

[499] Peter D. Sasieni and Patrick Royston. General interest section: Dotplots. Applied Statistics, 45(2):219–234, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bassett:1996:GIS

[500] E. E. Bassett, S. P. Brooks, and B. J. T. Morgan. General interest section: MINITAB multivariate macros. Applied Statistics, 45(2):235–245, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Raab:1996:LE

[501] G. M. Raab, P. Dellaportas, and A. F. M. Smith. Letter to the editors. Applied Statistics, 45(2):247–251, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crook:1996:SAA

[502] J. F. Crook and I. J. Good. Statistical algorithms: Algorithm AS 301: The computation of a Bayes factor against independence in contingency tables. Applied Statistics, 45(2):255–265, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/301.

Shi:1996:SAA

[503] Ning-Zhong Shi and Zhi Geng. Statistical algorithms: Algorithm AS 302: Multiple isotonic regression. Applied Statistics, 45(2):266–273, 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/302.

Kim:1996:SAR

[504] Minbo Kim. Statistical algorithms: Remark AS R96: A remark on Algorithm AS 279: Computing p-values for the generalized Durbin-Watson statistic and residual autocorrelations in regression. Applied Statistics, 45(2): 273–274, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/279. See [296].

Anonymous:1996:BMPb

[505] Anonymous. Back matter. Applied Statistics, 45(2):[4], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1996:FMPb

[506] Anonymous. Front matter. Applied Statistics, 45(2):[5], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Satten:1996:MCM

[507] Glen A. Satten and Ira M. Longini, Jr.

Markov chains with measurement error: Estimating the ‘true’ course of a marker of the progression of human immunodeficiency virus disease. Applied Statistics, 45(3):275–309, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). With discussion.

Bisgaard:1996:CAP

[508] Søren Bisgaard. A comparative analysis of the performance of Taguchi’s linear graphs for the design of two-level fractional factorials. Applied Statistics, 45 (3):311–322, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1996:KTT

[509] Martin J. Crowder. Keep timing the tablets: statistical analysis of pill dissolution rates. Applied Statistics, 45 (3):323–334, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Raghunathan:1996:MIA

[510] Trivellore E. Raghunathan and David S.

Siscovick. A multiple-imputation analysis of a case-control study of the risk of primary cardiac arrest among pharmacologically treated hypertensives. Applied Statistics, 45(3):335–352, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Have:1996:MCA

[511] Thomas R. Ten Have. Marginal and conditional association models for analysing departures from exchangeability of tooth bone loss. Applied Statistics, 45(3):353–369, 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gange:1996:UBB

[512] Stephen J. Gange, Alvaro Munoz, Marc Saez, and Jordi Alonso. Use of the betabinomial distribution to model the effect of policy changes on appropriateness of hospital stays. Applied Statistics, 45 (3):371–382, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1996:LE

[513] Philip J. Smith, Theodore J. Thompson, Thomas R. Ten Have, and David H. Uttal. Letter to the editors. Applied Statistics, 45(3):383–384, 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyons:1996:SAA

[514] N. I. Lyons and K. Hutcheson. Statistical algorithms: Algorithm AS 303: Generation of ordered multinomial frequencies. Applied Statistics, 45(3): 387–393, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/303.

Richards:1996:SAA

[515] Larry E. Richards and John Byrd. Statistical algorithms: Algorithm AS 304: Fisher’s randomization test for two small independent samples. Applied Statistics, 45(3):394–398, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/304.

Preston:1996:CSD

[516] Ian Preston. Corrigendum: Sampling distributions of relative poverty statistics. Applied Statistics, 45(3):399, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [429].

Anonymous:1996:BMPc

[517] Anonymous. Back matter. Applied Statistics, 45(3):[3], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1996:FMPc

[518] Anonymous. Front matter. Applied Statistics, 45(3):[5], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nester:1996:ASC

[519] Marks R. Nester. An Applied Statistician’s creed. Applied Statistics, 45(4):401–410, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=035&iid=4&vid=
45.

Betensky:1996:ACM

[520] Rebecca A. Betensky and Alice S. Whittemore. An analysis of correlated multivariate binary data: Application to familial cancers of the ovary and breast. Applied Statistics, 45(4):411–429, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=036&iid=4&vid=
45.

Thall:1996:EGC

[521] Peter F. Thall, Derek Jacoby, and Stuart O. Zimmerman. Estimating genomic category probabilities from fluorescent in situ hybridization counts with misclassification. Applied Statistics, 45(4):431–446, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=037&iid=4&vid=
45.

Babcock:1996:EST

[522] Bruce A. Babcock, Alicia L. Carriquiry, and Hal S. Stern. Evaluation of soil test information in agricultural decision-making. Applied Statistics, 45(4):447–461, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=038&iid=4&vid=
45.

Coles:1996:BAE

[523] Stuart G. Coles and Jonathan A. Tawn. A Bayesian analysis of extreme rainfall data. Applied Statistics, 45(4):463–478, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=039&iid=4&vid=

Denis:1996:ACR

[524] Jean-Baptiste Denis and John C. Gower.

Asymptotic confidence regions for biadditive models: Interpreting genotypeenvironment interactions. Applied Statistics, 45(4):479–493, 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=040&iid=4&vid=
45.

Caroni:1996:LE

[525] Chrys Caroni, Colin Aitken, Wing K.

Fung, Kay I. Penny, P. W. Tsai, S. G. Gilmour, and R. Mead. Letters to the editors. Applied Statistics, 45(4): 495–503, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gan:1996:SAA

[526] F. F. Gan. Statistical algorithms: Algorithm AS 305: Average run lengths for cumulative sum control charts under linear trend. Applied Statistics, 45 (4):505–512, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/305.

Moser:1996:SAA

[527] Barry Kurt Moser and Melinda H. Mc-Cann. Statistical algorithms: Algorithm AS 306: The balanced incomplete block product. Applied Statistics, 45 (4):512–516, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/306.

Rousseeuw:1996:SAA

[528] Peter J. Rousseeuw and Ida Ruts. Statistical algorithms: Algorithm AS 307: Bivariate location depth. Applied Statistics, 45(4):516–526, 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/307.

Anonymous:1996:BMPd

[529] Anonymous. Back matter. Applied Statistics, 45(4):[3], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1996:FMPd

[530] Anonymous. Front matter. Applied Statistics, 45(4):[5], 1996. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1996:VIP

[531] Anonymous. Volume information. Applied Statistics, 45(4):[3], 1996. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lange:1997:NLF

[532] Nicholas Lange and Scott L. Zeger. Nonlinear Fourier time series analysis for human brain mapping by functional magnetic resonance imaging. Applied Statistics, 46(1):1–29, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=046&iid=1&vid=

46. With discussion and a reply by the authors.

Diggle:1997:SAR

[533] Peter J. Diggle and Ibrahim al Wasel.

Spectral analysis of replicated biomedical time series. Applied Statistics, 46(1):31–71, December 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=047&iid=1&vid=

46. With discussion and a reply by the authors.

Lunn:1997:MCM

[534] David J. Lunn and Leon J. Aarons.

Markov chain Monte Carlo techniques for studying interoccasion and intersubject variability: Application to pharmacokinetic data. Applied Statistics, 46(1):73–??, December 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=048&iid=1&vid=
46.

Rai:1997:DSM

[535] S. N.. Rai and D. E. Matthews. Discrete scale models for survival-sacrifice experiments. Applied Statistics, 46 (1):93–109, December 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=049&iid=1&vid=
46.

Ridout:1997:ECP

[536] M. S. Ridout and D. C. Harris. Estimating the concentration of propagules of a plant pathogen from soil dilution plate counts when the response is non-linear. Applied Statistics, 46(1):111–??, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=050&iid=1&vid=
46.

Smith:1997:SEA

[537] Richard L. Smith. Statistics for exceptional athletics records. Applied Statistics, 46(1):123–128, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=051&iid=1&vid=
46.

Turkkan:1997:AHP

[538] N. Turkkan and T. Pham-Gia. Algorithm AS 308: Highest posterior density credible region and minimum area confidence region: The bivariate case. Applied Statistics, 46(1): 131–??, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/308;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=053&iid=1&vid=
46.

Jaimez:1997:AEM

[539] R. Gutiérrez Jáimez, A. González Carmona, and F. Torres Ruiz. Algorithm AS 309: Estimation in multivariate lognormal diffusion processes with exogenous factors. Applied Statistics, 46 (1):140–146, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/309;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=054&iid=1&vid=
46.

Chattamvelli:1997:ACN

[540] R. Chattamvelli and F. Shanmugam.

Algorithm AS 310: Computing the non-central beta distribution function. Applied Statistics, 46(1):146– 156, December 1997. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/310;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=055&iid=1&vid=
46.

Mauricio:1997:AEL

[541] José Alberto Mauricio. Algorithm AS 311: The exact likelihood function of a vector autoregressive moving average model. Applied Statistics, 46 (1):157–??, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/311;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=056&iid=1&vid=
46.

Chan:1997:AAS

[542] Grace Chan and Andrew T. A. Wood.

Algorithm AS 312: An algorithm for simulating stationary Gaussian random fields. Applied Statistics, 46 (1):171–??, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/312;

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=057&iid=1&vid=
                                        46.

Garcia-Perez:1997:CAE

[543] Miguel A. García-Pérez. Correction to Algorithm AS 179: Enumeration of all permutations of multi-sets with fixed repetition numbers, by Miguel A. García-Pérez. Applied Statistics, 46(1):182, December 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

                                        //lib.stat.cmu.edu/apstat/179;
                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=058&iid=1&vid=
                                        46.

Anonymous:1997:ASIa

[544] Anonymous. Algorithms section: Introduction. Applied Statistics, 46(1):??, December 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=052&iid=1&vid=
                                        46.

Robinson:1997:SES

[545] Michael. E. Robinson and Jonathan A. Tawn. Statistics for extreme sea currents. Applied Statistics, 46 (2):183–??, 1997. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

                                        //www.blackwellpublishers.co.uk/
                                        asp/journal.asp?ref=0035-9254&src=
                                        ard&aid=059&iid=2&vid=46.

John:1997:CET

[546] J. A. John and E. R. Williams. The construction of efficient two-replicate row column designs for use in field trials. Applied Statistics, 46(2):207–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=060&iid=2&vid=
46.

Gibson:1997:MCM

[547] Gavin J. Gibson. Markov chain Monte Carlo methods for fitting spatiotemporal stochastic models in plant epidemiology. Applied Statistics, 46(2):215–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=061&iid=2&vid=
46.

Jacob:1997:FCH

[548] G. Jacob, F. H. C. Marriott, and P. A. Robbins. Fitting curves to human respiratory data. Applied Statistics, 46(2):235–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=062&iid=2&vid=
46.

Lindsey:1997:PMI

[549] J. K. Lindsey. Parametric multiplicative intensities models fitted to bus motor failure data. Applied Statistics, 46(2):245–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=063&iid=2&vid=

46. See corrigendum [578].

Marriott:1997:GTS

[550] J. M. Marriott and A. N. Pettitt. Graphical techniques for selecting explanatory variables for time series data. Applied Statistics, 46(2):253–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=064&iid=2&vid=
                                        46.

Dixon:1997:MAF

[551] Mark J. Dixon and Stuart G. Coles. Modelling association football scores and inefficiencies in the football betting market. Applied Statistics, 46(2):265–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=065&iid=2&vid=
                                        46.

Nelder:1997:LE

[552] J. A. Nelder and G. J. S. Ross. Letters to the editors. Applied Statistics, 46(2):281–286, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=066&iid=2&vid=
                                        46.

Donev:1997:AAC

[553] A. N. Donev. Algorithm AS 313: An algorithm for the construction of crossover trials. Applied Statistics, 46(2):288–295, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/313;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=068&iid=2&vid=
46.

Payne:1997:AIM

[554] R. W. Payne. Algorithm AS 314: Inversion of matrices with contents subject to modulo arithmetic. Applied Statistics, 46(2):295–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/314;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=069&iid=2&vid=
46.

Anonymous:1997:ASIb

[555] Anonymous. Algorithms section: Introduction. Applied Statistics, 46 (2):??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://www.

blackwellpublishers.co.uk/asp/journal.
asp?ref=0035-9254&src=ard&aid=067&
iid=2&vid=46.

Donnelly:1997:BSE

[556] C. A. Donnelly, S. M. Gore, R. N.

Curnow, and J. W. Wilesmith. The bovine spongiform encephalopathy maternal cohort study: Its purpose and findings. Applied Statistics, 46(3):299–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-

                                        9254&src=ard&aid=070&iid=3&vid=
                                        46.

Gore:1997:BSE

[557] Sheila M. Gore, Walter R. Gilks, and John W. Wilesmith. Bovine spongiform encephalopathy maternal cohort study — exploratory analysis. Applied Statistics, 46(3):305–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=071&iid=3&vid=
                                        46.

Donnelly:1997:ABS

[558] C. A. Donnelly, A. C. Ghani, N. M. Ferguson, J. W. Wilesmith, and R. M. Anderson. Analysis of the bovine spongiform encephalopathy maternal cohort study: Evidence for direct maternal transmission. Applied Statistics, 46(3):321–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=072&iid=3&vid=
                                        46.

Curnow:1997:ABS

[559] R. N. Curnow, A. Hodge, and J. W. Wilesmith. Analysis of the bovine spongiform encephalopathy maternal cohort study: The discordant case-control pairs. Applied Statistics, 46(3):345–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=073&iid=3&vid=
                                        46.

Goetghebeur:1997:ETC

[560] Els Goetghebeur and Krista Lapp.

The effect of treatment compliance in a placebo-controlled trial: Regression with unpaired data. Applied Statistics, 46(3):351–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=074&iid=3&vid=
46.

Wright:1997:NDE

[561] David Wright, Julian Stander, and Kypros Nicolaides. Nonparametric density estimation and discrimination from images of shapes. Applied Statistics, 46(3):365–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=075&iid=3&vid=
46.

Mendoza-Blanco:1997:AAS

[562] José R. Mendoza-Blanco and Xin M.

Tu. Algorithm AS 315: An algorithm for sampling the degrees of freedom in Bayesian analysis of linear regressions with t-distributed errors. Applied Statistics, 46(3):383–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/315;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=077&iid=3&vid=
46.

Moser:1997:ARM

[563] Barry Kurt Moser and Melinda H. Mc-Cann. Algorithm AS 316: A Robbins-Monro-based sequential procedure. Applied Statistics, 46(3):388–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/316;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=078&iid=3&vid=
46.

Thompson:1997:FMM

[564] T. J. Thompson, P. J. Smith, and J. P. Boyle. Finite mixture models with concomitant information: assessing diagnostic criteria for diabetes. Applied Statistics, 46(3):393–404, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=117&iid=3&vid=
46.

Agha:1997:AML

[565] M. Agha and D. S. Branker. Algorithm AS 317: Maximum likelihood estimation and goodness-of-fit tests for mixtures of distributions. Applied Statistics, 46(3):399–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/317;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=079&iid=3&vid=
46.

Tiwari:1997:AER

[566] Ram C. Tiwari and Jie Yang. Algorithm AS 318: An efficient recursive algorithm for computing the distribution function and non-centrality parameter of the non-central F -distribution. Applied Statistics, 46(3):408–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/318;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=080&iid=3&vid=
46.

Anonymous:1997:ASIc

[567] Anonymous. Algorithms section: Introduction. Applied Statistics, 46 (3):??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://www.

blackwellpublishers.co.uk/asp/journal.
asp?ref=0035-9254&src=ard&aid=076&
iid=3&vid=46.

Fitzmaurice:1997:DOL

[568] Garrett M. Fitzmaurice, Anthony F.

Heath, and D. R. Cox. Detecting overdispersion in large scale surveys: Application to a study of education and social class in Britain (with comments). Applied Statistics, 46(4):415–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=081&iid=4&vid=
46.

Volinsky:1997:BMA

[569] Chris T. Volinsky, David Madigan, Adrian E. Raftery, and Richard A. Kronmal. Bayesian model averaging in proportional hazard models: Assessing the risk of a stroke. Applied Statistics, 46(4):433–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-

                                        9254&src=ard&aid=082&iid=4&vid=
                                        46.

Olguin:1997:NLH

[570] Jorge Olguín and Tom Fearn. A new look at half-normal plots for assessing the significance of contrasts for unreplicated factorials. Applied Statistics, 46(4):449–462, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=083&iid=4&vid=
                                        46.

Lipsitz:1997:QRM

[571] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Geert Molenberghs, and Lue Ping Zhao. Quantile regression methods for longitudinal data with drop-outs: Application to CD4 cell counts of patients infected with the human immunodeficiency virus. Applied Statistics, 46(4):463–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=084&iid=4&vid=
                                        46.

Stephens:1997:CIE

[572] D. A. Stephens, A. F. M. Smith, and R. Moskovic. Charpy impact energy data: A Markov chain Monte Carlo analysis. Applied Statistics, 46(4):477–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=085&iid=4&vid=
                                        46.

Taplin:1997:SAP

[573] Ross H. Taplin. The statistical analysis of preference data. Applied Statistics, 46(4):493–512, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=086&iid=4&vid=
46.

Koval:1997:AVM

[574] John J. Koval. Algorithm AS 319: Variable metric function minimization. Applied Statistics, 46(4):515–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/319;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=088&iid=4&vid=
46.

Garcia-Perez:1997:ADI

[575] Miguel A. García-Pérez. Algorithm AS 320: Decomposing an integer N into all sets of J positive integer summands by simulation of dynamically varying nested DO loops. Applied Statistics, 46(4):522–533, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/320;
http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=089&iid=4&vid=
46.

Hilton:1997:AEH

[576] Joan F. Hilton and Lauren Gee. Algorithm AS 321: An exact Hollander bivariate symmetry test. Applied Statistics, 46(4):533–??, 1997. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

                                        //lib.stat.cmu.edu/apstat/321;
                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=090&iid=4&vid=
                                        46.

Gilks:1997:CAR

[577] W. R. Gilks, R. M. Neal, N. G. Best, and K. K. C. Tan. Corrigendum: Adaptive rejection Metropolis sampling. Applied Statistics, 46(4):541–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=091&iid=4&vid=

46. See [466].

Lindsey:1997:CPM

[578] J. K. Lindsey. Corrigendum: Parametric multiplicative intensities models fitted to bus motor failure data. Applied Statistics, 46(4):542–??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=092&iid=4&vid=

46. See [549].

Anonymous:1997:AIV

[579] Anonymous. Author index: Vol. 46, 1997. Applied Statistics, 46(4):??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=093&iid=4&vid=
                                        46.

Anonymous:1997:AS

[580] Anonymous. Algorithms section: Introduction. Applied Statistics, 46 (4):??, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://www.

blackwellpublishers.co.uk/asp/journal.
asp?ref=0035-9254&src=ard&aid=087&
iid=4&vid=46.

Bartholomew:1998:SUC

[581] D. J. Bartholomew. Scaling unobservable constructs in social science. Applied Statistics, 47(1):1–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=094&iid=1&vid=
47.

Ehrlich:1998:SRA

[582] Willa K. Ehrlich, Vijayan N. Nair, M. S. Alam, W. H. Chen, and Mark Engel. Software reliability assessment using accelerated testing methods. Applied Statistics, 47(1):15–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=095&iid=1&vid=
47.

Aslett:1998:COS

[583] Robert Aslett, Robert J. Buck, Steven G. Duvall, Jerome Sacks, and William J. Welch. Circuit optimization via sequential computer experiments: Design of an output buffer. Applied Statistics, 47(1):31–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=096&iid=1&vid=
                                        47.

Kvam:1998:BFR

[584] Paul H. Kvam. The binomial failure rate mixture model for common cause failure data from the nuclear industry. Applied Statistics, 47(1):49–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=097&iid=1&vid=
                                        47.

Curnow:1998:LGI

[585] R. N. Curnow, A. P. Morris, and J. C. Whittaker. Locating genes involved in human diseases. Applied Statistics, 47(1):63–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=098&iid=1&vid=
                                        47.

Ventura:1998:SIE

[586] V. Ventura, A. C. Davison, and S. J. Boniface. Statistical inference for the effect of magnetic brain stimulation on a motoneurone. Applied Statistics, 47(1):77–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=099&iid=1&vid=
                                        47.

Donner:1998:SAD

[587] Allan Donner. Some aspects of the design and analysis of cluster randomization trials. Applied Statistics, 47(1):95–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=100&iid=1&vid=
47.

Spiegelhalter:1998:BGM

[588] David J. Spiegelhalter. Bayesian graphical modelling: A case-study in monitoring health outcomes. Applied Statistics, 47(1):115–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=101&iid=1&vid=
47.

Zhou:1998:CAT

[589] Xiao-Hua Zhou. Comparing accuracies of two screening tests in a twophase study for dementia. Applied Statistics, 47(1):135–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=102&iid=1&vid=
47.

Lindsey:1998:AHS

[590] J. K. Lindsey and P. M. E. Altham. Analysis of the human sex ratio by using overdispersion models. Applied Statistics, 47(1):149–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=103&iid=1&vid=
47.

Copas:1998:OGR

[591] John Copas and Peter Marshall. The offender group reconviction scale: A statistical reconviction score for use by probation officers. Applied Statistics, 47(1):159–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=104&iid=1&vid=
47.

Hamilton:1998:CNL

[592] David C. Hamilton and O. Knop. Combining non-linear regressions that have unequal error variances and some parameters in common. Applied Statistics, 47(2):173–185, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=105&iid=2&vid=
47.

Lin:1998:NNL

[593] T. C. Lin and M. Pourahmadi. Nonparametric and non-linear models and data mining in time series: a casestudy on the Canadian lynx data. Applied Statistics, 47(2):187–201, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=106&iid=2&vid=
47.

Steiner:1998:GDE

[594] Stefan H. Steiner. Grouped data exponentially weighted moving average control charts. Applied Statistics, 47(2):203–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=107&iid=2&vid=
47.

Longford:1998:CDT

[595] N. T. Longford. Count data and treatment heterogeneity in 2 × 2 crossover trials. Applied Statistics, 47(2):217–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=108&iid=2&vid=
47.

Faraway:1998:TSF

[596] Julian Faraway and Chris Chatfield.

Time series forecasting with neural networks: A comparative study using the airline data. Applied Statistics, 47(2):231–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=109&iid=2&vid=
47.

Ekholm:1998:MCO

[597] Anders Ekholm and Chris Skinner. The muscatine children’s obesity data reanalysed using pattern mixture models. Applied Statistics, 47(2):251–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=110&iid=2&vid=
                                        47.

Ridout:1998:MMM

[598] Martin S. Ridout, Malcolm J. Faddy, Graeme Jones, and Colin A. M. Campbell. Markov models of mite movements. Applied Statistics, 47(2):265–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=111&iid=2&vid=
                                        47.

Dellaportas:1998:BCN

[599] Petros Dellaportas. Bayesian classification of Neolithic tools. Applied Statistics, 47(2):279–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=112&iid=2&vid=

47. See corrigendum [619].

Diggle:1998:MBG

[600] P. J. Diggle, J. A. Tawn, and R. A. Moyeed. Model-based geostatistics. Applied Statistics, 47(3):299–350, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=113&iid=3&vid=

47. With discussion and a reply by the authors.

Maritz:1998:EGC

[601] J. S. Maritz, C. J. Lombard, and C. H. Morrell. Exact group comparisons using irregular longitudinal data. Applied Statistics, 47(3):351–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=114&iid=3&vid=
47.

Hurn:1998:CFM

[602] M. Hurn. Confocal fluorescence microscopy of leaf cells: An application of Bayesian image analysis. Applied Statistics, 47(3):361–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=115&iid=3&vid=
47.

Lewis:1998:CDP

[603] S. M. Lewis and K. G. Russell. Crossover designs in the presence of carry-over effects from two factors. Applied Statistics, 47(3):379–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=116&iid=3&vid=
47.

Thompson:1998:FMM

[604] T. J. Thompson, P. J. Smith, and J. P. Boyle. Finite mixture models with concomitant information: Assessing diagnostic criteria for diabetes. Applied Statistics, 47(3):393–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bruun:1998:CAE

[605] J. T. Bruun and J. A. Tawn. Comparison of approaches for estimating the probability of coastal flooding. Applied Statistics, 47(3):405–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=118&iid=3&vid=
                                        47.

Troxel:1998:ALD

[606] A. B. Troxel, D. P. Harrington, and S. R. Lipsitz. Analysis of longitudinal data with non-ignorable nonmonotone missing values. Applied Statistics, 47(3):425–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=119&iid=3&vid=
                                        47.

Best:1998:NAO

[607] B. J. Best and J. C. W. Rayner. Nonparametric analysis of ordinal categorical response data with factorial structure. Applied Statistics, 47(3):439–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=120&iid=3&vid=
                                        47.

Bissell:1998:LEH

[608] A. F. Bissell. Letters to the Editors: Half-normal plots for unreplicated factorials. Applied Statistics, 47(3):447, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=121&iid=3&vid=
47.

Nelder:1998:LEI

[609] J. A. Nelder. Letters to the Editors:

The importance of marginality rules. Applied Statistics, 47(3):447–448, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=122&iid=3&vid=
47.

Angelis:1998:BPA

[610] D. De Angelis, W. R. Gilks, and N. E. Day. Bayesian projection of the acquired immune deficiency syndrome epidemic (with discussion). Applied Statistics, 47(4):449–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=123&iid=4&vid=
47.

Pawitan:1998:MIP

[611] Y. Pawitan. Mixed inverse problems arising in the estimation of calibration factors in positron emission tomography. Applied Statistics, 47(4):499–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=124&iid=4&vid=
47.

Dittrich:1998:MES

[612] R. Dittrich, R. Hatzinger, and W. Katzenbeisser. Modelling the effect of subject-specific covariates in paired comparison studies with an application to university rankings. Applied Statistics, 47(4):511–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=125&iid=4&vid=

47. See corrigendum [662].

Mertens:1998:EPC

[613] B. J. A. Mertens. Exact principal component influence measures applied to the analysis of spectroscopic data on rice. Applied Statistics, 47(4):527–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=126&iid=4&vid=
                                        47.

Christen:1998:SSR

[614] J. A. Christen and C. E. Buck. Sample selection in radiocarbon dating. Applied Statistics, 47(4):543–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=127&iid=4&vid=
                                        47.

Kelsall:1998:SVR

[615] J. E. Kelsall and P. J. Diggle. Spatial variation in risk of disease: A nonparametric binary regression approach. Applied Statistics, 47(4):559–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=128&iid=4&vid=
47.

Vounatsou:1998:BAT

[616] P. Vounatsou, T. Smith, and A. F. M.

Smith. Bayesian analysis of twocomponent mixture distributions applied to estimating malaria attributable fractions. Applied Statistics, 47 (4):575–??, 1998. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//www.blackwellpublishers.co.uk/
asp/journal.asp?ref=0035-9254&src=
ard&aid=129&iid=4&vid=47.

Bray:1998:AMC

[617] I. Bray and D. E. Wright. Application of Markov chain Monte Carlo methods to modelling birth prevalence of Down syndrome. Applied Statistics, 47(4):589–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=130&iid=4&vid=
47.

Hadgu:1998:BAL

[618] A. Hadgu and Y. Qu. A biomedical application of latent class models with random effects. Applied Statistics, 47(4):603–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=131&iid=4&vid=

Dellaportas:1998:CBC

[619] Patros Dellaportas. Corrigendum: Bayesian classification of Neolithic tools. Applied Statistics, 47(4):620–??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=133&iid=4&vid=

47. See [599].

Anonymous:1998:CV

[620] Anonymous. Contents of volume 47, 1998. Applied Statistics, 47(4):??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=134&iid=4&vid=
                                        47.

Anonymous:1998:IV

[621] Anonymous. Index to Volume 47, 1998. Applied Statistics, 47(4):??, 1998. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=132&iid=4&vid=
                                        47.

Allum:1999:EBE

[622] G. T. Allum, R. G. Aykroyd, and J. G. B. Haigh. Empirical Bayes estimation for archaeological stratigraphy. Applied Statistics, 48(1):1–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=135&iid=1&vid=
                                        48.

Hughes:1999:NHH

[623] J. P. Hughes, P. Guttorp, and S. P.

Charles. A non-homogeneous hidden Markov model for precipitation occurrence. Applied Statistics, 48(1):15–30, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//www.blackwellpublishers.co.uk/
asp/journal.asp?ref=0035-9254&src=
ard&aid=136&iid=1&vid=48.

Faddy:1999:SMI

[624] M. J. Faddy and J. S. Fenlon. Stochastic modelling of the invasion process of nematodes in fly larvae. Applied Statistics, 48(1):31–??, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=137&iid=1&vid=
48.

Oman:1999:CTM

[625] S. D. Oman, N. Meir, and N. Halm.

Comparing two measures of creatinine clearance: An application of errorsin-variables and bootstrap techniques. Applied Statistics, 48(1):39–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=138&iid=1&vid=
48.

Avery:1999:FMC

[626] P. J. Avery and D. A. Henderson. Fitting Markov chain models to discrete state series such as DNA sequences. Applied Statistics, 48(1):53–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=139&iid=1&vid=
48.

Rosenbaum:1999:UQA

[627] P. R. Rosenbaum. Using quantile averages in matched observational studies. Applied Statistics, 48(1):63–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=140&iid=1&vid=
48.

Weller:1999:PRS

[628] E. A. Weller, E. J. Feuer, C. M. Frey, and M. N. Wesley. Parametric relative survival regression using generalized linear models with application to Hodgkin’s lymphoma. Applied Statistics, 48(1):79–??, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=141&iid=1&vid=
48.

Gilmour:1999:CPF

[629] S. G. Gilmour and T. J. Ringrose. Controlling processes in food technology by simplifying the canonical form of fitted response surfaces. Applied Statistics, 48(1):91–101, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=142&iid=1&vid=
48.

Liu:1999:IHI

[630] X. Liu, C. Waternaux, and E. Petkova.

Influence of human immunodeficiency virus infection on neurological impairment: An analysis of longitudinal binary data with informative drop-out. Applied Statistics, 48(1):103–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=143&iid=1&vid=
48.

Raab:1999:ISB

[631] G. M. Raab and C. A. Donnelly. Information on sexual behaviour when some data are missing. Applied Statistics, 48(2):117–??, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=144&iid=2&vid=
48.

Dixon:1999:ENS

[632] Mark J. Dixon and Jonathan A. Tawn.

The effect of non-stationarity on extreme sea-level estimation. Applied Statistics, 48(2):135–151, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=145&iid=2&vid=
48.

GuttierrezToscano:1999:UCC

[633] P. Guttiérrez Toscano and F. H. C.

Marriott. Unsupervised classification of chemical compounds. Applied Statistics, 48(2):153–163, 1999.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=146&iid=2&vid=
48.

Copas:1999:ERS

[634] John Copas. The effectiveness of risk scores: The logit rank plot. Applied Statistics, 48(2):165–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=147&iid=2&vid=
48.

Ridout:1999:MVB

[635] Martin S. Ridout, Byron J. T. Morgan, and David R. Taylor. Modelling variability in the branching structure of strawberry inflorescences. Applied Statistics, 48(2):185–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=148&iid=2&vid=
48.

Webb:1999:BFA

[636] Andrew R. Webb and Philip N. Garner. A basis function approach to position estimation using microwave arrays. Applied Statistics, 48(2):197–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=149&iid=2&vid=
48.

Souza:1999:STA

[637] K. M. A. de Souza, J. T. Kent, and K. V. Mardia. Stochastic templates for aquaculture images and a parallel pattern detector. Applied Statistics, 48(2):211–??, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=150&iid=2&vid=
48.

Faddy:1999:MAD

[638] M. F. Faddy and M. C. Jones.

Modelling and analysis of data that exhibit temporal decay. Applied Statistics, 48(2):229–??, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=151&iid=2&vid=
48.

Stanghellini:1999:DVC

[639] E. Stanghellini, K. J. McConway, and D. J. Hand. A discrete variable chain graph for applicants for credit. Applied Statistics, 48(2):239–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=152&iid=2&vid=
48.

Langford:1999:MMG

[640] Ian H. Langford, Alistair H. Leyland, Jon Rasbash, and Harvey Goldstein. Multilevel modelling of the geographical distributions of diseases. Applied Statistics, 48(2):253–??, 1999.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=153&iid=2&vid=
48.

Verbyla:1999:ADE

[641] Arũnas P. Verbyla, Brian R. Cullis, Michael G. Kenward, and Sue J. Welham. The analysis of designed experiments and longitudinal data by using smoothing splines (with discussion). Applied Statistics, 48(3):269–311, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=154&iid=3&vid=
48.

Sauerbrei:1999:URM

[642] Willi Sauerbrei. The use of resampling methods to simplify regression models in medical statistics. Applied Statistics, 48(3):313–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=155&iid=3&vid=
48.

Kelsall:1999:FDL

[643] Julia E. Kelsall, Scott L. Zeger, and Jonathan M. Samet. Frequency domain log-linear models; air pollution and mortality. Applied Statistics, 48(3):331–??, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-

9254&src=ard&aid=156&iid=3&vid=
48.

Sanso:1999:VRD

[644] Bruno Sansó and Lelys Guenni.

Venezuelan rainfall data analysed by using a Bayesian space-time model. Applied Statistics, 48(3):345–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=157&iid=3&vid=
48.

Verbeke:1999:EDE

[645] Geert Verbeke and Emmanuel Lesaffre. The effect of drop-out on the efficiency of longitudinal experiments. Applied Statistics, 48(3):363–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=158&iid=3&vid=
48.

Glickman:1999:PEL

[646] Mark E. Glickman. Parameter Estimation in Large Dynamic Paired Comparison Experiments. Applied Statistics, 48(3):377–394, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=159&iid=3&vid=
48.

Durham:1999:CTS

[647] L. Kathryn Durham, M. Elizabeth Halloran, Ira M. Longini, Jr., and Amita K. Manatunga. Comparison of two smoothing methods for exploring waning vaccine effects. Applied Statistics, 48(3):395–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=160&iid=3&vid=
48.

Lambert:1999:AFR

[648] Phillipe Lambert and J. K. Lindsey. Analysing financial returns by using regression models based on nonsymmetric stable distributions. Applied Statistics, 48(3):409–??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=161&iid=3&vid=
48.

Hadgu:1999:A

[649] A. Hadgu and Y. Qu. Acknowledgement. Applied Statistics, 48(3):425, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=162&iid=3&vid=
48.

Avery:1999:DCS

[650] P. J. Avery and D. A. Henderson. Detecting a changed segment in DNA sequences. Applied Statistics, 48(4):53–61, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=167&iid=4&vid=
48.

Denison:1999:MTS

[651] D. G. T. Denison, A. T. Walden, A. Balogh, and R. J. Forsyth. Multitaper testing of spectral lines and the detection of the solar rotation frequency and its harmonics. Applied Statistics, 48(4):427–439, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=163&iid=4&vid=
48.

Trinca:1999:DVD

[652] L. A. Trinca and S. G. Gilmour. Difference variance dispersion graphs for comparing response surface designs with applications in food technology. Applied Statistics, 48(4):441–455, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=164&iid=4&vid=
48.

Breslow:1999:DAT

[653] N. E. Breslow and N. Chatterjee. Design and analysis of two-phase studies with binary outcome applied to Wilms tumour prognosis. Applied Statistics, 48(4):457–468, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=165&iid=4&vid=
48.

Barao:1999:EAS

[654] M. I. Barão and J. A. Tawn. Extremal analysis of short series with outliers: sea-levels and athletics records. Applied Statistics, 48(4):469–487, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=166&iid=4&vid=
48.

Gower:1999:ADS

[655] J. C. Gower and W. J. Krzanowski. Analysis of distance for structured multivariate data and extensions to multivariate analysis of variance. Applied Statistics, 48(4):505–519, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=168&iid=4&vid=
48.

Khuri:1999:UQP

[656] A. I. Khuri, J. M. Harrison, and J. A. Cornell. Using quantile plots of the prediction variance for comparing designs for a constrained mixture region: an application involving a fertilizer experiment. Applied Statistics, 48(4):521–532, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=169&iid=4&vid=
48.

Heagerty:1999:SER

[657] P. J. Heagerty and M. S. Pepe. Semiparametric estimation of regression quantiles with application to standardizing weight for height and age in US children. Applied Statistics, 48(4):533–551, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=170&iid=4&vid=
48.

Lindsey:1999:UCO

[658] J. K. Lindsey. On the use of corrections for overdispersion. Applied Statistics, 48(4):553–561, 1999. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=171&iid=4&vid=
48.

Smith:1999:ANI

[659] P. W. F. Smith, C. J. Skinner, and P. S. Clarke. Allowing for nonignorable non-response in the analysis of voting intention data. Applied Statistics, 48(4):563–577, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=172&iid=4&vid=
48.

Anonymous:1999:CV

[660] Anonymous. Contents of volume 48, 1999. Applied Statistics, 48(4):??, 1999. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=174&iid=4&vid=
48.

Anonymous:1999:IAV

[661] Anonymous. Index of authors, volume 48, 1999. Applied Statistics, 48(4):??, 1999. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

                                        //www.blackwellpublishers.co.uk/
                                        asp/journal.asp?ref=0035-9254&src=
                                        ard&aid=173&iid=4&vid=48.

Dittrich:2001:CME

[662] R. Dittrich, R. Hatzinger, and W. Katzenbeisser. Corrigendum: Modelling the effect of subject-specific covariates in paired comparison studies with an application to university rankings. Applied Statistics, 50(2):??, 2001. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [612].